### v1

# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

Abstract.
Greenwashing is an effort to mislead the public about the carbon transition of an entity, such as a state or company. We provide a comprehensive survey of the scientific literature addressing natural language processing methods to identify potentially misleading climate-related corporate communications, indicative of greenwashing. We break the detection of greenwashing into intermediate tasks, and review the state-of-the-art approaches for each of them. We discuss datasets, methods, and results, as well as limitations and open challenges. We also provide an overview of how far the field has come as a whole, and point out future research directions.
natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
1. Introduction
Since the first report of the Intergovernmental Panel on Climate Change in 1990  (Change, 1990), world leaders and citizens have increasingly recognized the significant adverse impacts of human activities on the environment, particularly on climate change (Peter Jackson, 2007). This increased awareness has translated into guidelines, laws, and investments, such as the European Green Deal (Publications Office of the European Union, 2019) or the Inflation Reduction Act in the US (The White House, 2023). Many companies have used the financial incentives offered by states, and the guidelines and legislation to make significant steps towards sustainability (Time, 2024). At the same time, this growing attention also generated an advertising opportunity for companies that aim to promote themselves as environmentally aware and responsible. Indeed, some companies have been found to deliberately manipulate their data and statistics to appear more environment-friendly. The Diesel Scandal around the Volkswagen car company is a prominent example (US Environmental Protection Agency, 2024). However, such cases are not the norm. More commonly, companies avoid outright data manipulation but present themselves in a misleadingly positive light regarding their environmental impact – a practice called greenwashing.
In addition to their broader communication strategies, most companies disseminate information about their environmental initiatives through dedicated sustainability 10-K and annual reports. In addition, corporations also leverage various media channels and traditional communication platforms to convey their environmental credentials. As a result, the corpus of documents used in textual analysis is diverse and heterogeneous.
In such documents, greenwashing can appear in multiple forms. The Green Claims Directive of the European Union (US Environmental Protection Agency, 2023) identifies four types of greenwashing:
  1. (1)
    using unreliable sustainability labels (e.g., H&M and Decathlon have used labels such as eco design and conscious and were prompted by the Netherlands Authority for Consumers and Markets to stop the practice in 2022 as the retailers did not clearly describe why the products received those labels (Netherlands Authority for Consumers and Markets (2022), ACM))
  2. (2)
    presenting legal requirements for the product as its distinctive features (e.g., McDonald’s advertises its reduction of plastic waste by using reusable cutlery (McDonald’s France, 2024), while this is a legal obligation since January 2023 in France (LegiFrance, 2020))
  3. (3)
    making green claims about the entire product when the claim applies only to a part/aspect of the product (e.g., Kohl’s and Walmart were fined by the U.S. Federal Trade Commission for the misleading claim that their products were made of bamboo, a sustainable material, when they were made out of rayon, which is derived from bamboo via a highly toxic process that removes any environmental benefits of bamboo (Lesley Fair, Federal Trade Commission, 2022)), and finally
  4. (4)
    making environmental claims for which the company cannot provide evidence (e.g., Keurig was fined by Canada’s Competition Bureau for misleading claims about the recycling process of its single-use coffee pods (Competition Bureau Canada, 2022)).
Greenwashing is a widely observed practice rather than a marginal occurrence: the Australian Competition & Consumer Commission performed a sweep analysis on 247 businesses in October 2022 and found that $57\%$ of them made misleading claims about their environmental performance (Australian Competition and Consumer Commission, 2023). The International Consumer Protection Enforcement Network, too, found in November 2020 that $40\%$ of around 500 websites of businesses covering different sectors contain misleading green claims (UK Competition and Markets Authority, 2021).
Addressing greenwashing is not only in the broader public interest but also in the company’s best interest, as the practice is often subject to legal sanctions prohibiting misleading or deceptive conduct, with businesses facing fines or lawsuits as a consequence (Mary Foley, Forbes, 2024). In addition, political awareness and public awareness of the issue via traditional and social media (Thomas Wagner, Bonpote, 2021) can harm the reputation of a company involved in the practice. These factors make greenwashing a risk factor that can affect investments and that should be monitored at the company and portfolio levels.
The question thus arises to what extent greenwashing can be detected automatically. Such automated detection could help investors, journalists, prosecutors, and activists spot potentially illicit behavior, but it could also help companies avoid unintentionally misleading communication. In this paper, we provide a comprehensive survey of natural language processing (NLP) methods to identify potentially misleading climate-related corporate communication that can be indicative of greenwashing. We have identified 61 works that propose ways to check one or more aspects of greenwashing in text. We cover the intermediary steps involved in a natural language processing (NLP) analysis of the environmental disclosure of a company, ranging from detecting environmental or climate-related text to greenwashing detection. We structure these works in a unified categorization and discuss tasks, datasets, methods, and results for each of them. Unlike other surveys on the topic (Moodaley and Telukdarie, 2023b; Bernini et al., 2023), which compile and discuss findings reported in the scientific literature, our approach dives into the actual methodologies of these works.
Our main contributions and findings are:
  • We provide the first survey on NLP-based detection of greenwashing that describes and systemizes the methods used in the literature. Our survey focuses on tasks and datasets but also contextualizes the performance of the models and provides future research directions.
  • We find no datasets of positive and negative examples of greenwashing. Hence, it is not possible to train or evaluate models on the task of identifying greenwashing.
  • While there are no datasets explicitly labeled for greenwashing, many works have studied intermediary tasks, such as detecting misleading claims, vague language, or environmental reporting inconsistencies. Researchers have used these characteristics to develop theoretical definitions of greenwashing and proposed measurable indicators to identify potential cases. Although not a direct substitute for labeled greenwashing datasets, these indicators provide tools for analyzing corporate disclosures and assessing the likelihood of greenwashing.
  • Open challenges remain: While many models have been proposed for greenwashing-related tasks, it is difficult to compare approaches and solutions as many studies do not report measures of uncertainty (e.g., confidence interval, multiple runs, hyper-parameters tuning), nor do they report simple baselines that help assess the difficulty of datasets such as random baseline or TF-IDF baselines. Even the intermediary tasks thus remain active areas of research.
Our survey is structured as follows. We start with a discussion of the related work on greenwashing surveys (Section 2), followed by the legal context around climate actions and the definition of greenwashing (Section 3). We then survey intermediate tasks that have been studied in the literature (Section 4): identifying climate-related statements; identifying topics aligned with frameworks such as the Task Force on Climate-related Financial Disclosures (TCFD) or Environmental, Social, and Governance (ESG) factors; assessing the presence of green claims; and examining sentiment, argumentation quality, deceptive practices, and stance to discern the tone and authenticity of a message. We then review approaches that aim to identify greenwashing as a whole (Section 5). We conclude by discussing challenges and opportunities in greenwashing detection (Sections 6 and 7). We note that we have an Appendix section that will be part of the supplemental online-only material.
2. Related Work
Moodaley and Telukdarie (2023b) analyze the literature on greenwashing, sustainability reporting, and the intersection with research in Artificial Intelligence (AI) and Machine Learning (ML). The authors provide an overview of how the literature has changed over time by examining the number of publications and the different keywords covered by such publications. Unlike our work, this survey does not describe the actual approaches of the publications. Bernini et al. (2023) review the definitions of greenwashing in business, management, and accounting. They conclude that there is no agreed-upon definition of greenwashing, except that it is defined as a gap between the company’s actions and its disclosure. In contrast to our study, the review provides only a brief overview of the techniques for detecting greenwashing. While our review focuses on the semantic and linguistic aspects of greenwashing identification, other works focus on quantitative approaches that can serve the same objective (Dao et al., 2024).
Other works are less directly related to our survey. For example, a growing community is focused on using AI to tackle climate change [^1] See for example https://www.climatechange.ai/. Rolnick et al. (2022) review how machine learning has been used and could be used to tackle climate change. They explore different sectors of the economy, such as electricity, transportation, industry, and agriculture, among others. The survey discusses climate investments and the effects of climate change on finance, but not greenwashing. Alonso et al. (2023) provide a survey on ML approaches for climate finance, also known in the literature as green finance or carbon finance. Climate finance is a subfield in economics dedicated to providing tools from financial economics to tackle climate change. Finally, several works study the impact of AI on climate change (Hershcovich et al., 2022; Kaack et al., 2022; Rohde et al., 2023; Verdecchia et al., 2023). Our work, in contrast, surveys approaches that use AI to detect greenwashing.
3. Preliminaries
3.1. Legal Context
In this section, we provide the legal context for the definition of greenwashing. We first present international and national laws, followed by how they are applied in the corporate context.
International and Nation Level Legislation and Guidelines on Climate Change Mitigation
The first international summit on the effect of humans on the environment was the United Nations Conference on the Human Environment, held in 1972 in Sweden (United Nations, European Commission, 2024). The Intergovernmental Panel on Climate Change (IPCC) was created in 1988 and released its first report in 1990  (Change, 1990). It was only in 1992 that the first treaty between countries was signed, the United Nations Framework Convention on Climate Change (UNFCCC) (United Nations, European Commission, 2024). The treaty only encouraged countries to reduce their emissions, and in 1997, the Kyoto Protocol set commitments for the countries to follow (United Nations Framework Convention on Climate Change, 2011). The 36 countries that participated in the Protocol reduced their emissions in 2008–2012 by a large margin in respect to the levels in 1990; however, the global emissions increased by $32\%$ in the same period. Since the creation of UNFCCC, the nations, i.e., the parties, have met yearly at the “Conference of the Parties” (COP).
The three main directions of climate change policy are reducing greenhouse emissions, promoting renewable energy, and improving energy efficiency. One major milestone was the introduction in 2005 of the European Union Emissions Trading System, the world’s first large-scale emissions trading scheme. The trading system limits the amount of CO2 emitted by European industries and covers $46\%$ of the EU’s CO2 emissions. It is estimated that the trading system reduced CO2 emissions in the EU by $3.8\%$ between 2008 and 2016 (Bayer and Aklin, 2020). At COP21, in 2015, 194 nations plus the European Union signed the Paris Agreement, a treaty by which the nations commit to keep the rise in global surface temperature below 2 °C (3.6 °F) above pre-industrial levels. To achieve this goal, each country sends a national climate action plan every five years, and the parties assess the collective progress made towards achieving the climate goals. The first such assessment took place at COP28, where it was established that an important direction to achieve the long-term goals of countries was to transition away from fossil fuels to renewable energy. In 2021, the European Climate Law was adopted: the EU commits to reducing its emissions by at least $55\%$ by 2030 with respect to the 1990 levels and becoming climate neutral by 2050 (Council of the EU Press release, 2021). Similarly, countries such as Canada (Government of Canada, 2024), Taiwan (The Republic of China Ministry of Environment(2023), Taiwan), South Korea (The Korean Law Information Center, 2024), and Australia (Government of New South Wales, 2023), among others, aim to achieve carbon neutrality by 2050 and have passed laws setting this goal.
Corporate Level Laws and Guidelines
Laws and regulations at international and national levels have repercussions on companies that must comply or risk fines. However, assessing if a company has taken the necessary steps is not trivial, given that some laws refer to how a company will operate in the future, for example, by polluting less. Even without laws, investors can be concerned about environmental or social issues; hence, companies have used voluntary disclosures for a long time.
One of the first organizations to provide standards for reporting on climate change or social aspects is the Global Reporting Initiative (GRI), with its first guidelines published in 2000 (GRI, 2023). In 2015, the Financial Stability Board and the Group of 20 created the Task Force for Climate-related Financial Disclosure (TCFD) guidelines on disclosure in response to shortcomings of COP21, in particular the lack of standards climate-related disclosure (Wikipedia contributors, 2024). In 2021, the International Sustainability Standards Board (ISSB) was created to establish standards for climate-related disclosure, and starting in 2024, the standard released by this board will be applied worldwide (IFRS Foundation, 2024). ISSB standards were aligned with GRI disclosure standards to make them complementary and interoperable (Global Reporting Initiative, 2023), while the ISSB standards are expected to take over TCFD (IFRS Foundation, 2024). Differently from the voluntary reporting standards of GRI, TCFD, and ISSB, the Corporate sustainability reporting law in the EU required the creation of the European Sustainability Reporting Standards (ESRS), which are mandatory for companies subject to EU law (Directorate-General for Financial Stability, Financial Services and Capital Markets Union, 2023). ESRS has high interoperability with GRI and ISSB. GRI, TCFD, ESRS and ISSB fall under the umbrella term of environmental, social, and governance (ESG) guidelines for company disclosure. The European Union created its first law obliging companies to provide non-financial disclosure reporting in 2014, the Non-Financial Reporting Directive, which focused disclosure on environmental and social aspects. In 2023, the EU expanded this legislation via the Corporate Sustainability Reporting Law. Some European countries anticipated this legislation with their own, for example, the 2001 New Economic Regulations Act in France. Switzerland, which is not part of the EU, has imposed a mandatory TCFD disclosure for large public companies, banks, and insurance companies starting from January 2024 (The Swiss Federal Council, 2022), and similar laws exist in New Zealand (New Zealand Ministry of Environment, 2023). While not compulsory, initiatives like the Carbon Disclosure Project (CDP) have played a pivotal role by standardizing responses related to the climate disclosure of a company via structured questionnaires, thereby facilitating more systematic and comparable reporting.
3.2. Definition of Greenwashing
A widely cited and comprehensive definition, synthesizing those commonly found in the literature, is provided by the Oxford English Dictionary (OED, 2023).
Definition 0.
Greenwashing: To mislead the public (or to counter public or media concerns) by falsely representing a person, company, product, etc., as environmentally responsible.
While individuals, companies, or countries can all engage in greenwashing, we will focus on climate-related greenwashing by companies in this survey, with the following definition:
Definition 0.
Corporate climate-related greenwashing: To mislead the public into falsely representing the effort made by a company to achieve its carbon transition.
As climate-related disclosures face increasing regulation and greenwashing poses significant risks to a company’s reputation, some companies adopt a strategy of silence, avoiding discussions about their environmental impact. This deliberate lack of communication is known as greenhushing (Letzing, 2022). These definitions imply that greenwashing is a deliberate act; however, in many cases, it results from an error or miscommunication by companies genuinely trying to showcase their sustainability efforts, and that are trying to best follow disclosure standards.
As shown by Definitions 3.1 and 3.2, greenwashing is not defined by easily identifiable properties but as a general concept. Since the concept is so unspecific, researchers focused on components indicative of potentially misleading communications but easier to define. Because of this, in this review we are mentioning “greenwashing” explicitly, but also paraphrasing it as “misleading communications”, “misrepresentation of the company’s environmental impact/stance/performance”, or mentioning only components of it such as “cheap talk”, “selective disclosure/transparency”, “deceptive techniques”, “biased narrative”. They should all be understood in the context of climate-related misleading communications, as components associated or indicative of potential greenwashing even if they are not synonymous.
4. Intermediary tasks
4.1. Pretraining Models on Climate-Related Text
The first step in applying a language model for a given task is typically the pretraining process, which involves training the model on relevant corpora. Although general-purpose models such as BERT (Devlin et al., 2019) and LLaMA (Touvron et al., 2023) have demonstrated strong performance in various tasks, they are inherently limited by the knowledge and vocabulary present in their training corpora. This limitation can lead to suboptimal results when these models are applied to highly specialized domains with unique terminology and concepts. Hence, domain-specific models are usually trained on targeted datasets to improve their performance in niche areas.
Task Description
A language model is pretrained on climate-related text, usually on the classical tasks of masked token prediction and next sentence prediction. The goal is to produce a domain-specific pre-trained language model that will be subsequently fine-tuned to specific tasks. There are two approaches: (1) training a model with a domain-specific corpus from scratch, or (2) further-training [^2] Further-training refers to the process of training a pretrained model on its original task (e.g., next-token prediction or masked-token prediction) using additional domain-specific data to specialize its knowledge for that domain. a generalist pretrained model on a domain-specific corpus.
Annotated Datasets
There have been several efforts to create a climate-related corpus. Webersinke et al. (2022) gathered climate-related news articles, research abstracts, and corporate reports. Vaghefi et al. (2022) introduced Deep Climate Change, a dataset composed of abstracts of articles from climate scientists, and built a corpus specific to climate research texts. Schimanski et al. (2023b) introduced a dataset that focuses on text related to Environment Social and Governance (ESG). Similarly, Mehra et al. (2022) built a dataset using text from the Knowledge Hub of Accounting for Sustainability for an ESG domain-specific corpus. More recently, Thulke et al. (2024) proposed a dataset comprising news, publications (abstracts and articles), books, patents, the English Wikipedia, policy and finance-related texts, Environmental Protection Agency documents, and ESG, and IPCC reports. Mullappilly et al. (2023) proposed a climate-specific multilingual dataset. Yu et al. (2024) published a pretraining dataset constructed with annual and sustainable reports from EU banks. Unfortunately, to the best of our knowledge, only Yu et al. (2024)’s ClimateBUG-data dataset is publicly available.
Supervised Solutions
Based on the above datasets, the models ClimateBERT, ClimateGPT-2, climateBUG-LM and EnvironmentalBERT, ESGBERT were proposed (Webersinke et al., 2022; Vaghefi et al., 2022; Yu et al., 2024; Schimanski et al., 2023b; Mehra et al., 2022). More recently, generative domain-specific models such as Llama-2 for ClimateGPT (Thulke et al., 2024) or Vicuna for Arabic Mini-ClimateGPT (Mullappilly et al., 2023) have also been proposed.
Performance of Models
The domain-specific models drastically improve the performance on domain-specific masked-language modeling (Webersinke et al., 2022; Yu et al., 2024) and next token prediction (Vaghefi et al., 2022). They were also evaluated on domain-specific downstream tasks. These tasks are either based on pre-existing datasets such as ClimateFEVER (Leippold and Diggelmann, 2020) or introduced by the authors, as in ClimateBERT’s climate detection (Webersinke et al., 2022). At this stage, we evaluate whether fine-tuning enhances downstream performance, with a detailed analysis of the tasks presented in subsequent sections.
Cl.BERT DisRoBERTa NB
Climatext (Spokoyny et al., 2023; Leippold and Varini, 2020) 85.14 86.06 83.39
Climate Detection (Webersinke et al., 2022) 99.1±$\pm 1$ 98.6±$\pm 0.8$
Sentiment (Webersinke et al., 2022) 83.8±$\pm 3.6$ 82.5±$\pm 4.6$
Net-zero/Reduction (Schimanski et al., 2023a) 96.2±$\pm 0.4$ 94.4±$\pm 0.6$
Cl.BERT DisRoBERTa TF-IDF
TCFD classification (Sampson et al., 2022) 0.852 0.819 0.867
Cl.BERT DisRoBERTa SVM
SciDCC (Spokoyny et al., 2023; Mishra and Mittal, 2021) 52.97 51.13 48.02
ClimaTOPIC (Spokoyny et al., 2023) 64.24 63.61 58.34
Cl.FEVER (claim) (Xiang and Fujii, 2023; Leippold and Diggelmann, 2020) 76.8 72
Cl.FEVER (claim) (Webersinke et al., 2022; Leippold and Diggelmann, 2020) 75.7±$\pm 4.4$ 74.8±$\pm 3.6$
Cl.FEVER (evid.) (Spokoyny et al., 2023; Leippold and Diggelmann, 2020) 61.54 61.54
Cl.Stance (Spokoyny et al., 2023; Vaid et al., 2022) 52.84 52.51 42.92
ClimateEng (Spokoyny et al., 2023; Vaid et al., 2022) 71.83 72.33 51.81
ClimaINS (Spokoyny et al., 2023) 84.80 84.38 86.00
ClimaBENCH (Spokoyny et al., 2023) 69.44 69.27
Nature (Schimanski et al., 2024a) 94.11 94.03
Cl.BERT SVM+ELMo SVM+BoW
Commitment&Actions (Bingler et al., 2023) 81 79 76
Specificity (Bingler et al., 2023) 77 76 75
Cl.BERT DisBERT SVM
Env. Claim (Stammbach et al., 2023) 83.8 83.7 70.9
Cl.BERT DisBERT LSTM
DARE sentiment (Xiang and Fujii, 2023) 87.4 89.9 88.2
(a) F1-scores of climateBERT when compared to a similarly sized models. ROC-AUC for TCFD classification.
RoBERTa EnvRoBERTa DisRoBERTa EnvDisRoB.
Environment (Schimanski et al., 2023b) 92.35±$\pm 2.29$ 93.19±$\pm 1.65$ 90.97±$\pm 2$ $92.35\pm 1.65$
Social (Schimanski et al., 2023b) 89.87±$\pm 1.35$ 91.90±$\pm 1.79$ 90.59±$\pm 1.03$ $91.24\pm 1.86$
Governance (Schimanski et al., 2023b) 77.03±$\pm 1.82$ 78.48±$\pm 2.62$ 76.65±$\pm 2.39$ 78.86±$\pm 1.59$
EnvDisRoB. ClimateBERT RoBERTa DisRoBERTa
Water (Schimanski et al., 2024a) 94.47±$\pm 1.37$ 95.10±$\pm 1.13$ 94.55±$\pm 0.86$ $94.98\pm 1.16$
Forest (Schimanski et al., 2024a) 95.37±$\pm 0.92$ 95.34±$\pm 0.94$ 94.78±$\pm 0.48$ 95.29±$\pm 0.65$
Biodiversity (Schimanski et al., 2024a) 92.76±$\pm 1.01$ 92.49±$\pm 1.03$ 92.46±$\pm 1.54$ 92.29±$\pm 1.23$
Nature (Schimanski et al., 2024a) 94.19±$\pm 0.81$ 93.50±$\pm 0.64$ 93.97±$\pm 0.26$ 93.55±$\pm 0.72$
(b) F1-scores of multiple models compared to EnvironmentalBERT (based on RoBERTa and distilRoBERTa).
GPT-2 climateGPT-2
ClimateFEVER (Vaghefi et al., 2022) 62 72
LLama-2-Chat (7B) ClimateGPT (7B)
Multiple Benchmarks (Thulke et al., 2024) 71.4 77.1
(c) F1-scores of ClimateGPT-2 and custom performances (average Accuracy and F1-score over multiple benchmarks) of ClimateGPT compared to similarly sized models.
Table 1. Performance of domain-specific models on domain-specific tasks. The figure displayed in the tables are the values reported by authors in the corresponding studies; with the following abbreviations: CL for Climate, Dis for Distil, EnvDisRob for EnvDistilRoBERTa, Evid. for evidences. Each row reports the performances of models fine-tuned in the same experimental setting. Detailed performances are reported in Appendix C.
As shown in Table 1(c), both ClimateGPT-2 and ClimateGPT significantly outperform the baseline model in their designated tasks. These models benefit from additional training on climate-related data, highlighting the advantages of domain adaptation. However, when trained from scratch, Thulke et al. (2024) report substantially lower performance, underscoring the importance of large-scale pretraining data. For smaller models (climateBERT, climateBUG-LM and EnvironmentalBERT), the domain-adaptation also enhanced the performance on domain-specific tasks, though the significance of this improvement is less pronounced. As detailed in Table 1(a), out of the 15 evaluations of fine-tuned domain-specific climateBERT (Webersinke et al., 2022) on climate-related tasks compared to the fine-tuned base model (distilRoBERTa), 4 reported an improvement compared to the base model (¿1%), 8 reported a marginal improvement (¡1%) and 3 indicate no measurable difference (¡=0%). When compared to DistilBERT in the last two tasks from Table 1(a), climateBERT exhibits a slight improvement in one case, while performing comparably in the other. Additionally, climateBERT achieves performance similar to the simple baselines in certain instances (Table 1(a)). For example, in Specificity Classification (Bingler et al., 2023), climateBERT outperforms an SVM with bag-of-word by 2%, and on the TCFD classification (Sampson et al., 2022) climateBERT performs slightly below the TF-IDF baseline model. Likewise, climateBUG-LM (Yu et al., 2024) demonstrates a slight improvement of less than 1% over BERT and FinBERT on a climate-related detection task. As shown in Table 1(c), the domain-adapted models shows a small improvement in the performance on the Environmental task, by 1.38% for distilRoBERTa and 0.84% for RoBERTa.
Insights
Domain-specific language models, particularly those developed for environmental and climate-related texts, offer an interesting avenue for improving model performance through enhanced knowledge within specialized domains. Theoretically, such models should outperform general models by leveraging their tailored vocabulary and contextual understanding. However, our findings suggest that the improvements provided by the domain-specific models on the proposed domain-specific downstream tasks are limited. Indeed, we found that fine-tuning general models on domain-specific datasets consistently yields competitive performance. However, these findings do not undermine the significance of existing works that have been fundamental in demonstrating the potential of domain-specific models, particularly in generating domain-specific text (MLM and next token prediction). Moreover, ClimateGPT (Thulke et al., 2024) demonstrated significant improvement for the domain-adapted Llama-2, showing that domain-adaptation is highly relevant. Future research could focus on designing tasks where a lack of domain understanding—whether in terms of vocabulary or knowledge—significantly hinders performance.
4.2. Climate-Related Topic Detection
The first step toward detecting climate-related greenwashing is identifying text addressing climate-related topics.
Task Description
Given an input sentence or a paragraph, output a binary label, “climate-related” or “not climate-related”.
Dataset Input Labels Positive Label Details
ClimateBug-data (Yu et al., 2024) sentences from Banks’ reports relevant/ irrelevant Climate change and sustainability (including ESG, SDGs related to the environment, recycling and more)
ClimateBERT’s climate detection (Bingler et al., 2023) paragraphs from reports 1/0 Climate policy, climate change or an environmental topic
Climatext (Wikipedia, 10-K, claims) (Leippold and Varini, 2020) sentences from Wikipedia, 10-Ks or web scraping 1/0 Directly related to climate-change
Climatext (wiki-doc) (Leippold and Varini, 2020) sentences from a Wikipedia page 1/0 Extracted from a Wikipedia page related to climate-change
Sustainable signals’s reviews (Lin et al., 2023b) online product reviews (user comments) relevant/ irrelevant Contains terms related to sustainability
Table 2. Label definitions of the datasets related to climate change and sustainability topic detection.
Unsupervised Solutions
The earliest work on identifying corporate climate-related text was conducted by Doran and Quinn (2009), where the authors collected a large corpus of 10-K filings from 1995 to 2008, and filtered climate-related reports using hand-selected keywords. While Doran and Quinn (2009) focused on identifying climate-related text within company reports to assess how businesses addressed climate change, other early efforts took a different approach by analyzing the external impacts of media attention toward climate change on companies. For instance, Engle et al. (2020) developed the WSJ Climate Change News Index based on identifying climate-related news articles in media sources using climate-change vocabulary frequency. This index served as a tool to measure the influence of climate-change attention in the media on companies. Gorovaia and Makrominas (2025) proposed measuring the environmental content of CSR (corporate social responsibility) reports using a lexicon-based approach, with the goal of comparing CSR reports of environmental violators to companies with clean records.
Annotated Datasets
More recent studies have introduced several annotated datasets aimed at climate-related corporate text classification. One of the most prominent is ClimaText (Leippold and Varini, 2020), a large dataset from diverse sources such as Wikipedia, U.S. Security and Exchange Commission 10-K filings, and web-scraped content. The dataset is designed to cover a broad range of topics and document types to help assess climate-related discussions across different domains. It is divided into multiple subsets: wiki-doc, Wikipedia, 10-k, claims. Climatext (Wiki-doc) is annotated automatically, while Climatext (Wikipedia, 10-k and claims) is annotated by humans. In contrast, ClimateBERT’s authors (Webersinke et al., 2022) provide a smaller, more focused dataset consisting of paragraphs from companies’ annual and sustainability reports. This dataset was created as a downstream task for evaluating the ClimateBERT model, focusing on corporate disclosures. The dataset was later extended by Bingler et al. (2023), increasing its size and refining the annotation to capture both specificity and sentiment related to corporate climate discourse (see Sections 4.44.6). Another large dataset, ClimateBUG-data (Yu et al., 2024), also focuses on corporate communications but in EU banks. Beyond corporate disclosures, Sustainable Signals (Lin et al., 2023b) introduces a dataset of product reviews, each classified based on its relevance to sustainability. In addition to focusing on consumer perspectives through reviews, this dataset also examines sustainability aspects in product descriptions, providing a broader understanding of how sustainability is communicated in the context of consumer products.
Labels and Guidelines
As shown in Table 2, although these datasets address similar tasks, they differ significantly in their label definitions. The scope of the labels varies, ranging from a narrow focus on climate change to broader topics such as sustainability and environmental impact. While the task of the Climatext (wiki-doc) dataset is to detect whether a sentence is related to climate change or not, the weak label actually means that the sentence originates from a Wikipedia page related to climate change. Which is not the same task, as it focuses on the source of the statement rather than the actual content or its relevance to climate-related topics. However, the weakly labeled dataset serves as a useful filtering mechanism, helping to identify sentences that are potentially interesting. It provides an initial pool of data that can be refined with human annotations. Human annotations ensure that the labels align with the task’s goals, something weak labels alone cannot achieve.
Supervised Solutions
The solutions proposed to tackle that task range from keyword-based models (Leippold and Varini, 2020; Bingler et al., 2023) to fine-tuning BERT-like models (Leippold and Varini, 2020; Webersinke et al., 2022; Garrido-Merchán et al., 2023; Bingler et al., 2023; Yu et al., 2024) (see Table LABEL:tab:models in the appendix for details). The performance of the best models is above 90% (Garrido-Merchán et al., 2023; Bingler et al., 2023; Yu et al., 2024) (see Table 12 in the appendix for details).
Insights
The task of determining whether a statement is climate-related has been extensively studied, with numerous datasets available to support research in this area. State-of-the-art fine-tuned models now achieve near-perfect performance on these datasets, indicating that, under current conditions, the task is effectively solved. However, the labels used across the studies vary in scope. This highlights the inherent challenge of isolating climate-related discourse, given its close association with broader environmental and sustainability topics. For example, whether to include only explicit references to climate change or incorporate broader environmental and sustainability discussions, can profoundly influence the results and interpretations of such analyses. Moreover, relying only on weak labels might result in tasks that fail to capture the complexities of the subject.
4.3. Thematic Analysis
Once a text is known to be related to climate, one aims to know the exact topic of the text. Companies may prioritize specific categories while underreporting others, potentially signaling selective transparency practices such as greenwashing or greenhushing.
Task Description
Given an input sentence or a paragraph, output a subtopic related to climate change. This is a multiclass classification task that can be either supervised or unsupervised, depending on the availability of labeled data. Alternatively, it can be framed as a clustering task with the goal of discovering latent subtopic structures.
Dataset Input Labels Label details
TCFD rec. (Bingler et al., 2022) paragraphs from corporate annual reports TCFD 4 main categories: Metrics and Targets, Risk Management, Strategy, Governance and General (see appendix A for details)
TCFD (Sampson et al., 2022) paragraphs from regulatory and discretionary reports TCFD 11 recommendations: Metrics and Targets (board’s oversight, management’s role), Risk Management (identified risk, impact, resilience), … (see Appendix A for details)
FineBERT’s ESG (Huang et al., 2020) Sentences from 10-K Environmental, Social, Governance, General Environmental - e.g., climate change, natural capital, pollution and waste, and environmental opportunities
ESGBERT’s ESG (Schimanski et al., 2023b) Sentences from reports and corporate news Environment, Social, Governance and None Environmental criteria comprise a company’s energy use, waste management, pollution, […] as well as compliance with governmental regulations. Special areas of interest are climate change and environmental sustainability.
Multilingual ESG classification (Lee and Kim, 2023) Sentences from Korean reports (English and Korean) Environment, Social, Governance, Neutral, and Irrelevant Environmental factors include the reduction of hazardous substances, eco-friendly management, climate change, carbon emission, natural resources, […]
SciDCC (Mishra and Mittal, 2021) News articles (Title, Summary, Body) Environment, Geology, Animals, Ozone Layer, Climate, etc. Category in which the article was published (Automatic Label)
Transaction Ledger (Jain et al., 2023) Transaction ledger entry (description of transaction) Accounting Adjustments, Administration, Advertising, Benefits & Insurance, etc. Standardized commodity classes
ClimateEng (Vaid et al., 2022) Tweets posted during COP25 filtered by keywords (relevant to climate-change) Ocean/Water, Politics, Disaster, Agriculture/Forestry, General Sub-categories of climate-change
ESGBERT’s Nature (Schimanski et al., 2024a) Paragraphs from reports General, Nature, Biodiversity, Forest, Water Multi-label Nature-related topics
ClimaTOPIC (Spokoyny et al., 2023) CDP responses (short texts) Adaptation, Buildings, Climate Hazards, Emissions, Water, etc. Category of the question (Automatic Label)
Table 3. Summary of datasets, labels, scopes, and sources for climate-related text classification tasks.
4.3.1. Classifying Climate-Related Financial Disclosure (TCFD)
The Task Force on Climate-Related Financial Disclosures (TCFD) proposes four main categories of climate-related disclosure (Governance, Strategy, Risk Management, and Metrics and Targets) and 11 recommendations for climate-related disclosures (see Appendix A for details). Finding texts associated to these categories in company communication can help identify gaps and inconsistencies in reporting.
Unsupervised Solutions
Ding et al. (2023) introduced a score to quantify the extent of climate-related content in corporate disclosures, as well as 4 TCFD-category-based similarity scores. The score is based on sentence similarity with sentences from TCFD referential documents. They identified discrepancies that could signal greenwashing by comparing these scores with actual carbon emission data. Auzepy et al. (2023) examined reports from banks endorsing the TCFD recommendations. They proposed a fine-grained analysis of TCFD categories using a zero-shot entailment classification method on sentences.
Annotated Datasets
Although unsupervised methods can be employed for thematic analysis, recent studies have demonstrated the value of curated datasets for categorizing corporate disclosures. For instance, Bingler et al. (2022) leveraged report headings to annotate paragraphs into the four main TCFD categories. Similarly, Sampson et al. (2022) assessed the quality of climate-related disclosures using a dataset of 162k sentences, where each sentence was labeled according to the TCFD recommendations by experts through a question answering (QA) process.
4.3.2. Classifying Environmental, Social, and Governance Disclosure (ESG)
Beyond climate-specific reporting, Environmental, Social, and Governance (ESG) disclosures represent a broader set of non-financial metrics that evaluate a company’s sustainability performance. ESG reporting has become a global standard, encompassing climate-related initiatives and other dimensions of corporate responsibility. The standardized nature of ESG metrics makes them valuable for computational models that aim to detect inconsistencies in corporate claims, such as overstating environmental impacts through greenwashing. However, as noted in Berg et al. (2022)’s aggregated analysis of ESG ratings, discrepancies persist across different evaluation frameworks. These inconsistencies raise the question of whether artificial intelligence can be harnessed to improve the consistency and reliability of ESG assessments.
Unsupervised Solutions
Rouen et al. (2023) proposed an ML algorithm to describe the content of ESG reports. They used industry-topic dictionaries to compute the topic frequency using a TF-IDF-based algorithm. They used that classification to identify selective disclosure. Mehra et al. (2022) experimented with sentence similarity to extract sentences that were the most relevant to environmental factors, and (Bronzini et al., 2023) selected sentences related to ESG using INSTRUCTOR-xl (Su et al., 2023). They both subsequently used the extracted sentences to predict ESG scores.
Annotated Datasets
Multiple works (Huang et al., 2020; Schimanski et al., 2023b; Lee and Kim, 2023) proposed datasets for topic classification with the labels: Environment, Social, Governance and General. Huang et al. (2020) focused on reports while Schimanski et al. (2023b) added corporate news. Schimanski et al. (2023b) introduced a larger dataset for pre-training domain-specific models and an evaluation dataset of 2,000 sentences. While most ESG-related reports are typically available in English, some local companies provide them only in their native language. To partially address this, Lee and Kim (2023) proposed a multilingual dataset constructed from Korean corporate reports.
4.3.3. Other Topics
Although TCFD and ESG are established frameworks for analyzing corporate communications, many other topics may be of interest. These topics can follow other frameworks such as CDP, but the thematic analysis can also be unsupervised, effectively discovering the topics mentioned.
Unsupervised Solutions
Kang and Kim (2022) analyzed the content of sustainable reports with thematic analysis. It was conducted using sentence similarity between the reports and content from the SDGs website. Yu et al. (2024) examined the themes covered in the climate-related sections of bank reports by employing a dual approach: using hand-selected keywords to target anticipated topics and applying clustering techniques to group climate-related sentences and uncover underlying themes. Brié et al. (2022) leveraged ClimateBERT to extract climate-related paragraphs from corporate documents. They then applied a Structural Topic Model (STM) (Margaret E. Roberts and Airoldi, 2016) to identify thematic clusters within the data. This STM-based approach was also used by Fortes and Le Guenedal (2020) to identify influential topics for EUR/USD exchange rate. Their findings highlighted a significant discrepancy between the European Central Bank (ECB) and the Federal Reserve (FED), with the ECB attributing greater importance to climate change in the context of financial stability than the FED. Such thematic analysis can be used to identify selective disclosure.
Annotated Datasets
Spokoyny et al. (2023) introduced ClimaTOPIC, a topic classification dataset derived from responses to the CDP questionnaires, organized into 12 categories that align with the thematic structure of the CDP questions. This can be used to identify sections relevant to a particular CDP topic (e.g., Emissions, Adaptation, or Energy), enabling the extraction of structured information from unstructured documents. Schimanski et al. (2024a) presents a dataset focused on nature-related topics (such as Water, Forest, and Biodiversity) extracted from annual reports, sustainability reports, and earnings call transcripts (ESGBERT Nature). While less explicitly linked to climate change, these topics often reflect the consequences of environmental degradation associated with climate issues. While NLP is often used to analyze corporate report texts, it can also be used to analyze tables. For example, Jain et al. (2023) proposes estimating Scope 3 emissions  [^3] Scope 3 emissions refer to all indirect greenhouse gas (GHG) emissions that occur in a company’s value chain, excluding emissions from the company’s own operations (Scope 1) and its purchased electricity (Scope 2). based on corporate transactions (see Section 4.10 for details). The first step involves grouping expenses into standardized commodity classes, for which they introduced a dataset of 8K examples of corporate expenses categorized accordingly. As previously discussed, social media and news articles offer a broader perspective for cross-referencing corporate communications. For example, Vaid et al. (2022) introduced ClimateEng, a climate-related topic classification dataset composed of tweets grouped into five categories: Disaster, Ocean/Water, Agriculture/Forestry, Politics, and General. This dataset can be used to detect greenwashing by comparing the public’s concerns with the narrative presented in corporate sustainability reports. Similarly, Mishra and Mittal (2021) published the Science Daily Climate Change dataset (SciDCC), consisting of 11,000 news articles grouped into 20 climate-related categories. This dataset offers another avenue for identifying greenwashing, as companies’ communications can be cross-referenced with external news coverage, potentially highlighting contradictions or omissions in corporate reports.
4.3.4. Models and Conclusion
Supervised Solutions
The most common solution proposed for these tasks is fine-tuning a Transformer model (Huang et al., 2020; Schimanski et al., 2023b; Lee and Kim, 2023; Bingler et al., 2022; Sampson et al., 2022; Vaid et al., 2022; Spokoyny et al., 2023; Schimanski et al., 2024a; Jain et al., 2023). To contextualize the performance of Transformer models, some studies also report performances of classical baselines (Huang et al., 2020; Spokoyny et al., 2023; Schimanski et al., 2024a; Jain et al., 2023; Bingler et al., 2022; Sampson et al., 2022). Bingler et al. (2022) proposed a custom approach combining logistic regression with features from a fine-tuned language model, while Sampson et al. (2022) also evaluated clustering techniques and stacked models. The performance of the best-performing model is systematically high (above 80%), except on SciDCC and ClimaTOPIC. Classical approaches (SVM, NB) also reach good performances, under-performing by less than 20% on all datasets (except on Bingler et al. (2022)’s dataset). The TF-IDF baseline even outperformed fine-tuned Transformers on Sampson et al. (2022)’s dataset. This shows that those topics have distinguishable vocabularies.
Automatic Labels
The performance gap observed in Bingler et al. (2022) may be attributed to using paragraph-level inputs rather than sentence-level inputs aggregated for classification, as fine-tuned models also achieve around 20% precision at the paragraph level. The lower performances on SciDCC(Mishra and Mittal, 2021) and ClimaTOPIC(Spokoyny et al., 2023) are likely caused by the automation of the labeling process. The labels are therefore not designed to be labels. In SciDCC there are labels that are highly similar (e.g. Endangered Animals and Extinctions), or labels that include other labels (e.g. Environment, Climate, Pollution). As they are categories from a journal, the categories changed in time. For ClimaTOPIC, the labels are question categories, which are designed to group questions not to precisely identify them. Therefore a question about ”the emission of a building” might fit in either Emissions or Building, yet it is assigned only one label.
Insights
Topic classification is a well-established area of research in natural language processing (NLP) and has been extensively applied to climate-related topics. Fine-tuned Transformer models proposed in the literature consistently demonstrate near-perfect performance on human-annotated datasets, indicating that the task is largely solved. When reported, classical keyword baselines also perform relatively well, achieving significantly better than random performance and approximately 80% of the performance of fine-tuned models, suggesting the topics have distinct vocabularies. Future research could investigate topics that are lexically similar but distinct in their subject or focus. Topic detection serves as a valuable tool for structuring documents and has been utilized to analyze the extent to which companies address specific topics, enabling the detection of selective disclosure (Bingler et al., 2023, 2022). However, it is critical to ensure that labels are well-defined and data is accurately annotated. Lower performance observed on automatically generated labels may reflect challenges in predictability, potentially due to difficulties caused by the automatic annotation. These works might be revisited to conduct human annotations.
4.4. In-depth Disclosure: Climate Risk Classification
Climate change can bring both risks and opportunities for companies. Potential risks are, e.g., reputational risks (e.g. environmental controversies), regulatory risks (e.g. new regulations on emissions), and physical risks (e.g. droughts impacting production). Opportunities are, e.g., financial opportunities (e.g. benefiting from grants that aim to support less polluting industries), market opportunities (e.g. electric cars becoming more popular with environmentally conscious clients), etc. If a company systematically avoids discussing climate-related risks or disproportionately emphasizes opportunities, it creates a biased narrative that can serve as an indicator of greenwashing.
Task Description
Given an input sentence or a paragraph, output “opportunity” or “risk” label. Some works focus only on risks, classifying them into types of risks (e.g., physical risk, reputational risk, regulatory risk, or transition risk)
Dataset Input Labels
Ask BERT’s Climate Risk (Kölbel et al., 2021) Sentences from TCFD’s example reports and non-climate-related sentences Risk type: Transition risk, physical risk, and general risk (no guidelines)
Climate Risk (Friederich et al., 2021) Paragraphs from European companies annual reports Risk type:Acute, Chronic, Policy & legal, Tech & Market, Reputational, and Negative (no guidelines)
ClimateBERT’s Sentiment (Bingler et al., 2023) Paragraphs from companies’ annual reports Risk or threat that negatively impacts an entity of interest (negative sentiment); or Opportunity arising due to climate change (positive sentiment); Neutral otherwise.
Sentiment Analysis (Xiang and Fujii, 2023) paragraphs from academic texts on climate change and health published between 2013 and 2020 Risk (negative) if it discusses climate change causing public health issues, serious consequences, or worsening trends, greenwashing. Opportunity (positive): highlights potential benefits, positive actions, or research addressing gaps. Neutral otherwise.
Table 4. Summary of datasets, labels, scopes, and sources for climate-related Risk classification.
Unsupervised Solutions
Similarly to climate-related detection, climate risk classification has been tackled with keyword-based approaches. Li et al. (2020), Kheradmand et al. (2021) and Chou et al. (2023) proposed using dictionaries of words related to climate risk to identify paragraphs dealing with climate risk. Li et al. (2020) constructed risk measures based on the frequency of the terms in the risk dictionaries. Chou et al. (2023) analyzed the topics mentioned in conjunction with physical and transition risks. Sautner et al. (2023) proposed using one dictionary for risk and one for opportunity classification.
Annotated Datasets
Several studies have introduced specialized annotated datasets for risk/opportunity classification. For instance, Kölbel et al. (2021) created a dataset for climate-related risk classification with three categories: physical risk, transition risk, and general risk, using active learning for annotation. Similarly, Friederich et al. (2021) developed an annotated dataset for risk classification with five labels, covering acute physical risk, chronic physical risk, policy and legal risks, technology and market risks, and reputational transition risks. Bingler et al. (2023) released a dataset as part of the ClimateBERT downstream tasks, focusing on classifying paragraphs from corporate reports into three categories: opportunity, neutral, or risk. Extending beyond corporate disclosures, Xiang and Fujii (2023) compiled a climate-related risk/opportunity classification dataset of academic texts from the Web of Science and Scopus.
Supervised Solutions
The solutions proposed are similar to the other tasks: fine-tuned Transformer models (Kang and Kim, 2022; Kölbel et al., 2021; Friederich et al., 2021; Bingler et al., 2023; Webersinke et al., 2022; Xiang and Fujii, 2023) and keyword-based-features with simple models (Kölbel et al., 2021; Friederich et al., 2021; Bingler et al., 2023) (see Table LABEL:tab:models in the appendix for details). The notable exception is Xiang and Fujii (2023), who evaluated LSTM-based solutions alongside the Transformers models. Most studies (Kang and Kim, 2022; Kölbel et al., 2021; Bingler et al., 2023; Webersinke et al., 2022; Xiang and Fujii, 2023) reported high performances, above 80%. Kölbel et al. (2021) and Bingler et al. (2023) also reported good performances for the keyword-based baselines (72% and 84%). The performance reported by Friederich et al. (2021) on climate risk tasks highlights distinct challenges across subtasks. When tasked with identifying the specific type of risk in sentences already known to be about climate risks, word-based models outperformed fine-tuned Transformers. This suggests that each risk type had a distinct vocabulary. In contrast, when working with a dataset that reflects real-world proportions of risk and non-risk examples—where the data is heavily imbalanced—classical models struggled due to the overlap in vocabulary between general and climate risk-related text.
Difference between sentiment and risk
It is important to distinguish risk/opportunity classification from sentiment analysis. While risk/opportunity classification is inspired by sentiment analysis (Bingler et al., 2023), they play a complementary role. A statement such as “With proactive climate risk management, we are ready to tackle extreme weather disruptions, ensuring resilience” is classified as positive using a sentiment analysis model (Pérez et al., 2021), and as mentioning a risk using a risk/opportunity classification model (Bingler et al., 2023). Sentiment analysis focuses on the form of the statement, while risk is about the content. As the literature on sentiment analysis (Wankhade et al., 2022) is broad, the analysis of the tone of the statement can be done using existing models (Polignano et al., 2022; Kang and Kim, 2022) complementing approaches on risk classification.
Insights
The performance of fine-tuned models might indicate that the task of identifying texts talking about risks is solved. However, the work by Friederich et al. (2021) shows limitations when experimenting with heavily imbalanced datasets, which actually correspond to real-world settings. Future research could focus on such skewed settings.
4.5. Green Claim Detection
Identifying greenwashing involves more than just finding text related to climate change, as a company might simply state factual information about the climate without making any commitments on it. Hence, we now focus on detecting green claims, i.e., claims that a product, service, or corporate practice either contributes positively to environmental sustainability or is less harmful to the environment than alternatives (Stammbach et al., 2023; Woloszyn et al., 2021). The European Commission calls these claims “environmental claims”:
Dataset Input Positive Label description
Green Claims (Woloszyn et al., 2021) Marketing Tweets Environmental (or green) advertisements refer to all appeals that include ecological, environmental sustainability, or nature-friendly messages that target the needs and desires of environmentally concerned stakeholders.
Environmental Claims (Stammbach et al., 2023) Paragraph from reports Environmental claims refer to the practice of suggesting or otherwise creating the impression […] that a product or a service is environmentally friendly (i.e., it has a positive impact on the environment) or is less damaging to the environment than competing goods or services […] In our case, claims relate to products, services, or specific corporate environmental performance.
Table 5. Label definitions of the datasets on green claims detection.
Task Description
Given an input sentence or a paragraph, output a binary label, “green claim” or “not green claim”.
Annotated Datasets
The most notable approaches aiming at identifying climate-related claims introduced annotated datasets (Woloszyn et al., 2021; Stammbach et al., 2023). Woloszyn et al. (2021) proposed to focus on social media marketing through the detection of green claims in tweets. Stammbach et al. (2023) proposed a dataset for environmental claim detection in paragraphs from corporate reports.
Supervised Solutions
Both Stammbach et al. (2023) and Woloszyn et al. (2021) evaluated fine-tuned Transformer models on the datasets. Stammbach et al. (2023) also evaluated classical approaches and experimented with a model fine-tuned on general claim detection (Arslan et al., 2020) applied to their domain-specific task (see Table LABEL:tab:models in the Appendix for details). Woloszyn et al. (2021) and Stammbach et al. (2023) demonstrated good performances with fine-tuned Transformer (above 84%), yet they have very different Inter-Annotator Agreement (IAA), with Krippendorff’s $\alpha=0.8223$ for Woloszyn et al. (2021) and $\alpha=0.47$ for Stammbach et al. (2023). This might be explained by the difference in context: marketing tweets might be easier to understand for humans and more self-sufficient compared to paragraphs extracted from reports. Additionally, Woloszyn et al. (2021) experimented with adversarial attack, showing the sensibility to character-swap and word-swap. This shows that the models tend to rely heavily on particular words, showing only a superficial understanding. While it is difficult to generalize that conclusion given the small size of the dataset, it still indicates that BERT-like architecture can over-fit or rely on superficial cues instead of building an accurate representation of the paragraph. It is essential to assess model performance in challenging scenarios, where the presence of nonsensical or noisy inputs can reveal the fragility of model comprehension.
Insights
Green claims have been analyzed in company reporting (Stammbach et al., 2023) and social media communication (Woloszyn et al., 2021). Fine-tuned models can solve the task rather well, even if challenges remain:
  • Inter annotator agreement: The annotator agreement remains low for green claim detection in reports (Stammbach et al., 2023).
  • Integration with existing claim detection literature: While claim detection is a well-established field with extensive literature, studies on environmental claims do not fully connect with existing research on claim detection. In particular, they do not distinguish between Claims, Verifiable Claims (claims that can be checked), and Check-worthy Claims (claims that are interesting to verify) (Panchendrarajan and Zubiaga, 2024).
  • Evaluation Sensitivity and Real-World Robustness: The literature shows that fine-tuned models are sensitive to adversarial attacks, meaning small perturbations of the text influence greatly the performance of the classifier. As for all tasks, it is also important evaluate models robustness in real-world settings. Poor data quality might induce perturbations in the texts, reducing the performance of models.
4.6. Green Claim Characteristics
Once we have established that a sentence is climate-related (Section 4.2) and that it is a claim about the company (Section 4.5), we can endeavor to further classify the claim into fine-grained categories. Table 6 shows various characteristics of claims that have been studied.
Dataset Input Labels
Implicit/Explicit Green Claims (Woloszyn et al., 2021) Marketing Tweets Implicit green claims raise the same ecological and environmental concerns as explicit green claims (see definition in Section 4.5), but without showing any commitment from the company. If the tweet does not contain a green claim then No Claim.
Specificity (Bingler et al., 2023) Paragraph from reports A paragraph is Specific if it includes clear, tangible, and firm-specific details about events, goals, actions, or explanations that directly impact or clarify the firm’s operations, strategy, or objectives. Non-specific otherwise.
Commitments & Actions(Bingler et al., 2023) Paragraph from reports A paragraph is a commitment or an action if it contains targets for the future or actions already taken in the past.
Net Zero/Reduction (Schimanski et al., 2023a) Paragraph from Net Zero Tracker (Lang et al., 2023) The paragraph contains either a Net-Zero target, a Reduction target or no target (None)
Table 6. Label definitions the datasets related to characterization of green claims.
Task Description
Given an input sentence or a paragraph labeled as a green claim, output a more fine-grained characterization of the claim. This is a multi-label classification task; the labels can be about the form (e.g. specificity) or the substance (e.g. action, targets, facts).
Annotated Datasets
Existing works characterized both the content and the form of claims. On the content dimension, there are datasets for identifying if the claim is about an action or a commitment (Bingler et al., 2023; Schimanski et al., 2023a). Bingler et al. (2023) introduce a dataset for the identification of commitments and actions (Yes/No), and Schimanski et al. (2023a) released one for the identification of reduction targets (net zero, reduction, general). On the form dimension, Woloszyn et al. (2021) characterized claims as implicit or explicit, and Bingler et al. (2023) annotated their dataset on the specificity of claims (Specific/Not specific). Bingler et al. (2023) ultimately used the specificity and commitment/action characteristics to identify cheap talks related to climate disclosure.
Supervised Solutions
Woloszyn et al. (2021); Schimanski et al. (2023a) and Bingler et al. (2023) experimented with fine-tuned Transformers. Bingler et al. (2023) also evaluated classical baselines (SVM, NB), and Schimanski et al. (2023a) experimented with GPT-3.5-turbo (see Table LABEL:tab:models in the Appendix for details). All approaches reached good performances, in particular fine-tuned Transformers reaching performances above 80%. The exception remains the Specificity classification with a F1-score of 77% (close to the baselines). This low performance might be intrinsic to the task. Indeed, humans have disagreement when distinguishing specific and unspecific claims: Bingler et al. (2023) measured a low IAA on Specificity (Krippendorff’s $\alpha=0.1703$). This might indicate that the task is not well defined in the first place.
Insights
The high performances of fine-tuned models indicate that this type of task is solved. Furthermore, the demonstrated ability of GPT-3.5 as a zero-shot classifier (Schimanski et al., 2023a) highlights the potential of large language models to classify these characteristics without requiring extensive annotated datasets. These characteristics are particularly useful for identifying statements that may indicate greenwashing. For example, Bingler et al. (2023) utilized the proportion of non-specific commitments as a Cheap Talk Index demonstrating their practical applications. However, characteristics such as specificity are inherently ambiguous and subjective, as evidenced by the low inter-annotator agreement (IAA). Therefore, future research could focus on disambiguating what constitutes a specific statement from a non-specific one more objectively.
4.7. Green Stance Detection
Beyond making claims about the environmental impact of their products and processes, companies contribute to environmental discussions and can have an impact on regulatory frameworks through their communications and industry presence. It is thus helpful to understand the stance of an organization on the existence and gravity of climate change, on climate mitigation and adaptation efforts, and on climate-related regulations.
Dataset Input Labels Label details
ClimateFEVER (evidence)  (Leippold and Diggelmann, 2020) A claim and an evidence sentence from Wikipedia Support, Refutes, Not Enough Information Determines the relation between a claim and a single evidence sentence
LobbyMap (Stance) (Morio and Manning, 2023) Page from a company communications (report, press release, …) Strongly supporting, Supporting, No or mixed position, Not supporting, Opposing Given the policy and the page, classifies the stance
Global Warming Stance Detection (GWSD) (Luo et al., 2020) Sentences from news about global warming Stance of the evidence (Agree, Disagree, Neutral) toward the claim: Climate-Change is a serious concern.
ClimateStance (Vaid et al., 2022) Tweets posted during COP25 filtered by keywords (relevant to climate-change) Stance towards climate change prevention: Favor, Against, Ambiguous. (Stance used as a broad notion including sentiment, evaluation, appraisal, …)
Stance on Remediation Effort (Lai and Chen, 2023) Texts extracted from the TCFD sections of the financial reports The text indicates support for climate change remediation efforts or refutation of such efforts. (No guidelines)
Related subtask
ClimateFEVER (claim)  (Leippold and Diggelmann, 2020) A claim and multiple evidence sentences from Wikipedia Support, Refutes, Debated, Not Enough Information Determines if a claim is supported by a set of retrieved evidence sentences
LobbyMap (Page) (Morio and Manning, 2023) Page from a company communications (report, press release, …) 1/0 Contains a stance on a remediation policy
LobbyMap (Query) (Morio and Manning, 2023) GHG emission regulation, Renewable energy, Carbon tax, … Classifies the remediation policy
Table 7. Summary of datasets, labels, scopes, and sources for climate-related stance detection.
Task Description
Given two input sentences or paragraphs, one labeled as the claim and one as the evidence, predict the stance between the two: supports, refutes or neutral. Some studies fix the claim and only vary the evidence (e.g. the claim is always Climate change poses a severe threat), training a model to predict the stance of the evidence in respect to the fixed claim. Other studies train a model to predict the relation between varying claims/evidences.
Related subtask
The first subtask is collecting the evidences when the claims are already available (e.g. if they were collected manually). In order to build these datasets, research can rely on simple heuristics such as downloading all tweets published during the COP25 filtered with keywords (Vaid et al., 2022). However, other researchers performed a more elaborate procedure. Leippold and Diggelmann (2020) proposed a pipeline for collecting evidence. To retrieve relevant evidence from Wikipedia for a given claim, the pipeline involves three steps: document-level retrieval using entity-linking and BM25 to identify top articles, sentence-level retrieval using sentence embeddings trained on the FEVER dataset to extract relevant sentences, and sentence re-ranking using a pretrained ALBERT model to classify and rank evidence based on relevance. The second one encompasses the broader process: identifying claims and finding evidence, before predicting the relation. Wang et al. (2021) used a generic claim detection model trained on ClaimBuster (Arslan et al., 2020) for the claims and Google Search API to collect evidence. Finally, the last subtask is to train multiple models on each step: identifying evidences, identifying the claim and classifying the stance. This is the approach used by (Morio and Manning, 2023).
Annotated Datasets
As described previously, researchers built datasets of evidence related to one claim while focusing stance in tweets on seriousness of climate change such as Global Warming Stance Detection (GWSD) (Luo et al., 2020), ClimateStance (Vaid et al., 2022) stance on climate change of tweets posted during the COP25, or Lai and Chen (2023)’s dataset which focuses on the stance toward climate change remediation efforts. Another research initiative built a dataset of claim-evidence pairs, ClimateFEVER, and then trained a model on the stance classification (Leippold and Diggelmann, 2020; Wang et al., 2021). Finally, Morio and Manning (2023) proposed a dataset to assess corporate policy engagement built upon LobbyMap, which tackles the 3 steps: finding pages with evidence, identifying the claims targeted by that evidence, and classifying the stance.
Supervised Solutions
The solution proposed for to tackle stance classification are fine-tuned Transformer models (Vaid et al., 2022; Vaghefi et al., 2022; Xiang and Fujii, 2023; Morio and Manning, 2023; Webersinke et al., 2022; Wang et al., 2021; Spokoyny et al., 2023; Lai and Chen, 2023; Luo et al., 2020) and classical approaches (Spokoyny et al., 2023; Morio and Manning, 2023; Luo et al., 2020). Vaid et al. (2022) also experimented with FastText and Xiang and Fujii (2023) with LSTM-based models (See Table LABEL:tab:models in the Appendix for details). The performances are quite heterogeneous, Lai and Chen (2023) reaching F1-scores around 90% on classification of stance on remediation efforts, Luo et al. (2020) around 72% on GWSD, while performance on ClimateStance could not exceed 60%. Performances on ClimateFEVER are also quite low, however, when selecting only non-ambiguous examples, Xiang and Fujii (2023) could reach performances around 80%. Finally, performances on LobbyMap (Morio and Manning, 2023) are quite low (between 31% and 57.3% depending on the strictness of the metric). Overall, the performances show that the datasets are challenging (see Table 19 in the Appendix for details).
Exhaustivity
Datasets such as ClimateFEVER (Leippold and Diggelmann, 2020) and LobbyMap (Morio and Manning, 2023) rely on automated construction methods that may not ensure exhaustivity of evidence coverage. ClimateFEVER utilizes BM25 and Wikipedia for evidence retrieval, which could result in missing relevant information, particularly when compared to more advanced retrieval methods. LobbyMap relies on the LobbyMap.org website, which was not designed to be exhaustive. These limitations should be further investigated.
Insights
If an organization presents itself as environmentally friendly while simultaneously promoting climate-skeptic narratives or opposing climate remediation efforts, it probably is creating a misleading portrayal of its environmental stance. On the contrary, if they are aligned, it supports the authenticity of the organization’s efforts. Fortunately, stance detection has been extensively studied, yielding promising results; however, current performance levels leave room for improvement. Future research should focus on enhancing both methods and datasets to address these gaps. While existing datasets provide strong foundations, they may suffer from a lack of exhaustivity, particularly in evidence retrieval and coverage, which requires further investigation.
4.8. Question Answering
Question answering (QA) is a known task:
Task Description
Given an input question and a set of resources (paragraphs or documents), produce an answer to the question.
Dataset Input Labels
ClimaQA  (Spokoyny et al., 2023) The text from a response to one of the CDP questions; and one of the questions from the CDP questionnaire 1: the response answers this question 0: The response does not answer this question, but another one
ClimateQA (Luccioni et al., 2020) Sentence from reports and a question based on the TCFD recommendations 1: the sentence answers the question 0: The sentences does not answer the question
ClimaINS  (Spokoyny et al., 2023) The text from a response to one of the questions from the NAIC questionnaire MANAGE, RISK PLAN, MITIGATE, ENGAGE, ASSESS, RISKS The labels correspond the 8 questions asked in the NAIC questionnaires
Table 8. Summary of datasets, labels, scopes, and sources for climate-related QA datasets.
QA can be used for climate-specific applications such as structuring the information related to climate change from documents (Luccioni et al., 2020; Schimanski et al., 2023a), building chatbots with climate-related knowledge to make it more accessible to non-experts (Vaghefi et al., 2023; Rony et al., 2022), or helping identify potentially misleading information (Ni et al., 2023).
Solution
The QA process can be divided into two main steps (Kraus et al., 2023): the retrieval step and the answer generation step. The retrieval step involves locating the relevant information or answer to the user’s query in external documents. The answer generation step involves formulating a response to the user’s query from (1) the information retrieved from documents in the first step or (2) the pre-trained internal knowledge of a model.
Retriever
Retriever systems can be generic, using techniques such as sentence similarity, BM25, or document tags to filter documents and retrieve specific passages. Schimanski et al. (2024b) published ClimRetrieve, a benchmark for climate-related information retrieval in corporate reports. They found that simple embedding approaches are limited. Therefore, several works have specifically focused on improving the identification of answers within climate-related documents (Luccioni et al., 2020; Spokoyny et al., 2023). Luccioni et al. (2020) reformulated TCFD recommendations as questions and annotated reports to identify answers to the questions. The dataset includes questions paired with potential answers, each labeled to indicate whether the answer addresses the question or not. Based on the QA model trained on this dataset, they also proposed a tool called ClimateQA. Spokoyny et al. (2023) focused on questions from the CDP questionnaire and the NAIC Climate Risk Disclosure survey. As the responses to those questionnaires are publicly available, they constructed two datasets, ClimaINS and ClimaQA. ClimaQA is similar to ClimateQA (question/potential answer pairs). On the contrary, ClimaINS is a QA dataset framed as a classification task: it contains the responses from the NAIC survey, and each response is labeled as answering one of the 8 questions of the survey. Once a model has been trained on these datasets, it can be used to search for answers to the questions in other documents (effectively retrieving the information from an unstructured document). For ClimaINS, the authors experimented with multiple fine-tuned Transformers. As ClimaQA is framed as a retrieval task, the authors used BM25, sentence-BERT, and ClimateBERT to rank possible answers. Using the methodology evaluated on ClimaQA, they proposed a system to extract responses to CDP questions directly from unstructured reports and automatically fill the questionnaire (Spokoyny et al., 2023). Luccioni et al. (2020) trained and evaluated a RoBERTa-based model called ClimateQA on finding answers to TCFD-based questions. They all reached good performance but with room for improvement (see Table 21 in the Appendix for details), demonstrating the feasibility of the task but also the need for further research.
Answer generation
After the retrieval, the second step of question answering is to generate an answer for the question (given input resources or not). The generation can be simple such as finding a particular information in a paragraph. For example, Schimanski et al. (2023a) used a QA model (RoBERTa SQuaD v2 (Rajpurkar et al., 2016)) to extract the target year, the percentage of reduction, and the baseline year from the reduction target of a company from an input text that contains an emission reduction target. While it is possible to rely on generalist models, as they are proficient few-shot learners (Brown et al., 2020; Thulke et al., 2024), multiple studies proposed using domain-specific models. Climate Bot (Rony et al., 2022) built a dataset of climate-specific QA. Given a question and a document (scientific/news), the model should find the span answering the question. Mullappilly et al. (2023) proposed a dataset of question-answer pairs based on ClimaBench (Spokoyny et al., 2023) and CCMRC (Rony et al., 2022) to train a model to generate answers. Thulke et al. (2024) introduced a climate-specific corpus of prompt/completion pairs for Instruction Fine-Tuning created by experts and non-experts. (Rony et al., 2022) finetuned an ALBERT model on their climate-specific QA dataset. Mullappilly et al. (2023) trained a Vicuna-7B on their climate-specific answer generation dataset. And Thulke et al. (2024) trained LLama-2-based models (ClimateGPT) on their domain-specific prompt/completion dataset. They also evaluated their ClimateGPT model alongside multiple LLMs on climate-related benchmarks. They found that the domain-specific models outperformed generalist ones (see Table 21 in the Appendix for details).
Retrieval Augmented Generation
By combining the last two components (a retriever and a QA system), one can develop Retrieval-Augmented Generation (RAG) systems. Rony et al. (2022) provide the first example of a RAG system in climate-related tasks based on sentence-BERT to retrieve documents and AlBERT to identify the answer. More recently, Ni et al. (2023) proposed ChatReport, a methodology based on ChatGPT for analyzing corporate reports through the lens of TCFD questions using RAG. The reports, the answers to the TCFD questions, and a TCFD conformity assessment are chunked and stored in a vector store for easy retrieval. Each chunk and/or answer can be retrieved and injected into the prompt. They added in the prompt the notions of greenwashing and cheap talk to invite the model to provide a critical analysis of the retrieved answers. Similarly, Vaghefi et al. (2023) introduced ChatClimate, a RAG-based pipeline to augment GPT-4 with knowledge about IPCC reports. While the previous LLM-based approaches rely on closed-source models, Mullappilly et al. (2023) and Thulke et al. (2024) proposed RAG-based pipelines relying on open-source models. Rony et al. (2022) is the only study that evaluated their RAG pipeline using classical metrics (F1-score, BLEU, METEOR). Unfortunately, those metrics are not well aligned with human judgment for text generation (Chen et al., 2019). Therefore Vaghefi et al. (2023) and Mullappilly et al. (2023) relied on human or ChatGPT evaluations. They conclude that their approaches improve on non-domain-specific RAG systems.
Insights
The main application of Question-Answering (QA) systems is to analyze complex documents, such as corporate climate reports, that can be particularly useful for greenwashing detection. Modern LLMs achieve strong results in QA due to instruction-following fine-tuning, which improves adherence to specific instructions, and Retrieval-Augmented Generation (RAG), which bases its answers on retrieved documents. However, retrieval systems remain a performance bottleneck (Maekawa et al., 2024). Furthermore, answer generation is challenging to evaluate (Sai et al., 2022), often requiring human feedback or advanced model-based assessments.
4.9. Classification of Deceptive Techniques
In analyzing climate-related discourse, it can be useful to identify rhetorical strategies that might obscure or misrepresent an entity’s stance. For example, expressing support for climate remediation policies on one side (Morio and Manning, 2023), and promoting arguments downplaying the urgency of climate change on the other (Coan et al., 2021) could signal a lack of authenticity. Another example could be promoting a product through misleading rhetoric, by, for example, claiming that a product is better because it is natural, which would be an “appeal to nature” – a fallacious argument (Vaid et al., 2022; Jain et al., 2023). Detecting these deceptive techniques could help identify misleading communications.
Dataset Input Labels (Labels details)
LogicClimate (Jin et al., 2022) texts from climatefeedback.org Faulty Generalization, Ad Hominem, Ad Populum, False Causality, … Classifies fallacies (Multi-label)
Neutralization Techniques (Bhatia et al., 2021) paragraphs from other previous works on climate-change Denial of Responsibility, Denial of Injury, Denial of Victim, Condemnation of the Condemner, … Classifies neutralization techniques
Contrarian Claims (Coan et al., 2021) paragraphs from conservative think tank No Claim, Global Warming is not happening, Climate Solutions won’t work, Climate impacts are not bad, … Classifies arguments into super/sub-categories of climate science denier’s arguments
Table 9. Summary of datasets, labels, scopes, and sources for tasks related to deceptive techniques in climate-related context.
Task Description
The goal is to classify statements into argumentative categories: fallacies, types of arguments, or rhetorical techniques.
Unsupervised Solutions
Oppong-Tawiah and Webster (2023) frame greenwashing as fake news. They propose to tackle this identification through the form perspective. They developed a profile-deviation-based method to detect greenwashing in corporate tweets by comparing linguistic cues (e.g., quantity, specificity, complexity, diversity, hedging, affect, and vividness) to theoretically ideal profiles of truthful and deceptive communication. They compute a greenwashing score as Euclidean distances.
Annotated Datasets
An organization seeking to justify limited action might adopt a rhetoric that downplays the urgency of global warming or dismisses the impact of negligent behavior. These kinds of climate contrarian arguments mostly fall into a finite set of categories (e.g. “Solutions will not work”, or “Human influence is not demonstrated”). Therefore, Coan et al. (2021) constructed a taxonomy of such contrarian claims and published a large dataset annotated with such claims. Similarly, Bhatia et al. (2021) proposed a dataset to classify neutralization techniques, i.e., rationalizations that individuals use to justify deviant or unethical behavior (e.g. Denial of the victim). Finally, Jin et al. (2022) studied fallacious arguments and how they apply to climate change more broadly. They constructed LogicClimate, a climate-specific dataset of sentences annotated with fallacies, using articles from the Climate Feedback website.
Supervised Solutions
Fine-tuned Transformer architectures and classical approaches have been evaluated on each dataset (Jin et al., 2022; Bhatia et al., 2021; Coan et al., 2021). Jin et al. (2022) experimented with an ELECTRA model, fine-tuned to use the structure of the argument. The performance of fine-tuned approaches ranges from 58.77% to 79%. The models struggle with fallacy detection in LogicClimate, but perform well with generalist fallacy detection, contrarian claims, and neutralization techniques. Thulke et al. (2024) experimented with multiple zero-shot approaches using multiple LLM on the binary classification of contrarian claims using Coan et al. (2021)’s dataset. Their climateGPT-70B and LLama-2-Chat-70B models both reached an F1-score of 72.5% (see Table 23 in the Appendix for details).
Insights
The low performance of models on tasks such as fallacy detection and neutralization classification indicates the inherent complexity of these tasks. Fallacy detection, for one, is known to be inherently subjective (Helwe et al., 2024). The detection of neutralization techniques, too, appears to be subjective, as even human annotators achieve only a moderate performance level of 70% F1-score. Furthermore, both LogicClimate (Jin et al., 2022) and the neutralization dataset (Bhatia et al., 2021) are very small in size. To address these limitations, future research could explore several directions, including increasing the size of the datasets, defining the labels more precisely, or permitting multiple correct annotations (Helwe et al., 2024) to acknowledge viable disagreement. Additionally, the analyses on rhetorical techniques could be combined with other analyses (such as stance detection) to search for communication patterns.
4.10. Environmental Performance Prediction
Greenwashing can be interpreted as a misrepresentation of the company or product’s environmental performance. The environmental performance is often summarized by a quantitative metric such as the ESG score, the Finch score [^4] The Finch Score is a sustainability rating system designed to help consumers make eco-friendly choices by evaluating products on a scale from 1 to 10. See https://www.choosefinch.com/, or CO2 emissions. These scores are usually not directly mentioned in the company reports, but have to be inferred based on company communications:
Task Description
Given an input company report, output an environment-related quantitative value (such as the ESG score or the amount of Carbon Emission), even if that value is not mentioned in the report. In variants of this task, the goal is to predict not a value, but a range for that value, out of a set of possible ranges.
Unsupervised Solutions
Jain et al. (2023) proposed to predict the Scope 3 emissions from the list of financial transactions of the company. They first performed classification of all transaction descriptions to map them to their emission factors (emission per \$ spent), and then compute the emissions of each transaction. For ESG score prediction, Bronzini et al. (2023) proposes to use LLMs as few-shot learners to extract triples of ESG category, action, and company from sustainable reports, and to construct a graph representation with few-shot learning. They demonstrated the triplet generation using Alpaca, WizardLM, Flan-T5, and ChatGPT on a few examples. They could analyze disclosure choices and company similarities using the constructed graph. More importantly, they also used the graph to interpret ESG scores through the interpretability analysis of OLS predictions, effectively predicting the ESG score.
Annotated Datasets
Mehra et al. (2022) proposed to predict the changes in the ESG score instead of the actual value. They constructed their dataset by extracting the three sentences most relevant to the environment from financial reports and associated them with the E score [^5] ESG scores are usually aggregated scores along multiple dimensions. In this study, they are focusing on the ”Environment” part of the ESG score.’s change and direction of change. Instead of relying on scores, Clarkson et al. (2020) proposed to focus on good/bad CSR (corporate social responsibility) performers. Their approach evaluates CSR performance based on linguistic style rather than content, aiming to identify whether language patterns alone influence the perception of CSR quality. Sharma et al. (2024) introduced a multi-modal dataset of the tweets of companies to predict a company’s ESG unmanaged risk. Finally, focusing on products and not companies, Lin et al. (2023b) introduced a dataset of online product descriptions and reviews used to predict the Finch Score. All these scores can be used to identify companies and products that are actually environmentally friendly, helping users distinguish actual sustainability from greenwashing. However, they can also be used to find discrepancies between the communicated green-ness of a product or company (e.g. the percentage of climate-related texts in reports) and the actual green-ness (e.g. ESG score or quantity of emissions), which might indicate cases of greenwashing.
Supervised Solutions
Mehra et al. (2022) evaluated fine-tuned Transformers on ESG change prediction. Clarkson et al. (2020) experimented with hand-chosen features plugged into random forest and SVM on CSR performers prediction. Lin et al. (2023b) also experimented with traditional approaches (e.g. gradient boosting) and custom architectures based on Transformer models on Finch score prediction. Sharma et al. (2024) introduced baseline models, which are simple image and text embedding models (e.g., CLIP or sentence-BERT) with a classification head on ESG score prediction. A detailed list of models is reported in Table LABEL:tab:models in the Appendix. Bronzini et al. (2023); Clarkson et al. (2020); Sharma et al. (2024) did not report baselines, which makes the study difficult to analyze. (Mehra et al., 2022) reported good accuracy, demonstrating that a few sentences hold a significant amount of information to predict ESG score changes. However, there is definitely room for improvement, indicating that the information selected is not sufficient. Lin et al. (2023a) reported the performance of the baseline (average score) reaching already mean squared error (MSE) of 11.7% already quite low. Their Transformer-based approach reached MSE of 7.4% improving slightly on classical approaches (Gradient Boosting reaching MSE of 8.2%). Detailed performances can be found in Table 24 in the Appendix.
Insights
We expected the prediction of ESG and CSR metrics to be difficult and require an extensive understanding of a company. However, Bronzini et al. (2023); Lin et al. (2023b); Mehra et al. (2022) and Clarkson et al. (2020) showed that it is possible to build strong predictors relying only on textual elements. This can be explained because analysts reward transparency (Bronzini et al., 2023), so the quantity and complexity of disclosure have strong predictive power (Clarkson et al., 2020). Although existing studies have examined specific dimensions, future research should investigate the interaction between form (Clarkson et al., 2020) and content (Bronzini et al., 2023), as this could help uncover inconsistencies that may indicate greenwashing practices.
5. Greenwashing Detection
After having discussed different subtasks of greenwashing detection, we now come to the final, all-englobing task:
Task Description
Greenwashing detection is the task of predicting if a text contains greenwashing. This is a binary classification task. Greenwashing detection can also be done at an aggregated scale (such as by a yearly indicator).
Greenhushing and Selective Disclosure
There is a large body of work on disclosure (as described in Section 4.3.1). Quantifying climate-related disclosure can highlight greenhushing and selective disclosure. Bingler et al. (2022) identify a lack of disclosure in the Strategy and Metrics&Targets categories, which they describe as cherry-picking. They also highlight that merely announcing support for the TCFD does not lead to an increase in disclosure – on the contrary, it is a practice called cheap talk. Similarly, Auzepy et al. (2023) conducted a more fine-grained analysis of TCFD-related disclosure in the banking industry. They also highlighted large differences across TCFD categories. In particular, they found a low disclosure rate on the fossil-fuel industry-related topic despite large investments in that sector.
Climate Communication as an Image-Building Strategy
Ding et al. (2023) and Chou et al. (2023) found a correlation between climate-related disclosure and carbon emission: companies that emit more tend to disclose more climate-related information. Bingler et al. (2023) proposed a Cheap talk index based on claims specificity. They reached the same conclusion that larger emitters tend to disclose more, but also that negative news coverage is correlated with cheap talk. Polignano et al. (2022) showed that while disclosing more, reports mostly focus on positive disclosure. Kang and Kim (2022) proposed a sentiment ratio metric that highlights overly positive corporate reports. They identified periods of overly positive communications that followed negative environmental controversies; in other words, companies that tried to rebuild their image after a controversy. Gorovaia and Makrominas (2025) analyzed CSR reports of environmental violators and companies with clean records. They studied an ensemble of variables: quantification of environmental content, readability score, and sentiment analysis. They found that violators publish longer, more positive, and less readable CSR reports with more environment-related content. This suggests that companies use climate disclosure as a tool to mitigate controversies and/or to improve their image. Zhao et al. (2023) propose to use the discrepancies between internal disclosure and social media perception of a company to identify potential greenwashing.
The Role of Disclosure Style in Perceived Commitment
Clarkson et al. (2020) found that companies that use a more complex language in their CSR disclosure are associated with a higher CSR rating. Schimanski et al. (2023b) concluded that a higher quantity of ESG communication is associated with higher ESG rating. Rouen et al. (2023) also highlighted the relationship between disclosure quantity and complexity with the ESG score. This might indicate that the linguistic style is a good predictor of the substance of the discourse, or that analysts are rewarding companies that communicate on ESG-related issues.
ESG Score in Greenwashing Detection
Lee and Raschke (2023) proposed a greenwashing index based on the difference between the ESG score and the ESG score, weighted by the quantity of communication on each topic (E, S and G). Sharma et al. (2024) define two types of risks: managed risk (the company is addressing it), and unmanaged risk (the company is not currently addressing it but it could). Based on these, they propose an ESG unmanaged risk score as a measure for environmental performance, which can then be used as a signal for identifying greenwashing.
Contrasting Stances: A Method to Identify Greenwashing in Climate Communications
Morio and Manning (2023) proposed to evaluate a company’s stance towards climate change mitigation policies. They hypothesize that companies with a mixed stance might be engaging in greenwashing. The hypothesis was not tested and remains unconfirmed; still, such methodology helps understand the narratives of companies around climate change. Coan et al. (2021) were able to analyze the type of contrarian claims used by conservative think-tank (CTTs) websites and contrarian blogs. They showed a shift over time from climate change denying arguments to opposing climate change mitigation policies. Despite this shift, they found that the primary donors of these think tanks continue to support organizations promoting a climate-denier narrative aimed at discrediting scientific evidence and scientists. Discrepancies between the stance extracted from official communications (Morio and Manning, 2023) on the one hand, and the stance of third-party media organizations such as think tanks founded by the company (Coan et al., 2021) on the other, might indicate greenwashing.
Defining the linguistic features of greenwashing
Oppong-Tawiah and Webster (2023) we propose to identify greenwashing using the linguistic profile of tweets. They estimate the deceptiveness of the text using a keyword-based approach and linguistic indicators (e.g. word quantity, sentence quantity). They found a correlation between greenwashing and lower financial performances. A scoring system based solely on linguistic elements has significant limitations. For example, the following tweet gets a high greenwashing score: “Read about #[company’s] commitment to a #lowcarbon future http://[company website]”. However, it is merely an invitation to read a Web page.
Insights
Several indicators for greenwashing have been proposed: overly positive sentiment (Kang and Kim, 2022), the use of specific linguistic cues (Oppong-Tawiah and Webster, 2023), discrepancies between ESG scores and corporate disclosures (Lee and Raschke, 2023; Sharma et al., 2024), ambiguous stances (Morio and Manning, 2023), and inconsistencies between social media perception and official disclosures (Zhao et al., 2023). While these approaches laid the theoretical groundwork for understanding indicators of greenwashing, they have a significant limitation: they are not empirically evaluated (Bernini et al., 2023). Only a few studies, such as that by Kang and Kim (2022), have made attempts to validate their signals against actual cases, but even these efforts have not resulted in a comprehensive dataset of greenwashing examples. This gap between theory and practice highlights the necessity of developing datasets containing real-world instances of greenwashing. Without empirical evaluation, the prediction that a company engages in greenwashing is unfounded at best, and misleading or even defamatory at worst. However, building such a dataset comes with many challenges. First, one would have to identify suitable sources of documents that are openly accessible and likely to contain greenwashing. Second, one would have to overcome the inconspicuous and subjective nature of greenwashing itself. Finally, the publication of any such dataset exposes the authors to charges of defamation by the companies they accuse of greenwashing.
6. Open Challenges in Greenwashing Detection
We discuss open challenges in the evaluation (Section 6.1), model robustness (Section 6.2), and data sources (Section 6.3).
6.1. Evaluation Methodology
Ambiguous Labels
Many tasks in greenwashing detection are framed as classification tasks. However, in some cases, the labels are not sufficiently distinguishable. For example, specificity classification (Bingler et al., 2023) is the task of distinguishing specific claims from unspecific claims. It turns out that this task is itself too unspecific, as the low inter-annotator agreement shows. However, even for tasks that seem easy to distinguish, some edge cases remain ambiguous, e.g., the distinction between “climate-related” and “not climate-related”. The climate detection model climateBERT (Bingler et al., 2023) classifies the sentence “This year, we are producing more than 10,000 [bicycle/electric bicycle], demonstrating our continued growth and commitment to meeting market demand.” as “climate-related” for bicycle but not for electric bicycle. While electric bicycles are often associated with solutions to reduce emissions, the statement is not explicitly making the link. If we consider indirect links, then non-electric bicycles are also solutions. There lies the question of which product is inherently linked to climate-change, and which one requires contextualization.
When labels in classification tasks are not clearly distinguishable, the task is not well-defined. Consequently, it does not mean much if models perform well on such a task. It merely means that the model has learned whatever label distribution the dataset had. Work in this area should thus always first check human inter-annotator agreement on the task to make sure it is meaningful.
Generated Text Evaluation
Some tasks are not classification tasks, but text generation tasks. Evaluating generated text is a challenging task (Sai et al., 2022), and different techniques are being used: BLEU, METEOR, human-based evaluation, and LLM-based evaluation (Mullappilly et al., 2023; Rony et al., 2022; Vaghefi et al., 2023; Thulke et al., 2024). BLEU and METEOR are not well-suited for evaluating tasks such as question answering  (Chen et al., 2019), as they rely only on n-gram similarity. Human evaluation also has its shortcomings. First, it is limited to smaller sizes and variations in settings. Second, human evaluation of generated texts should be based on explicit guidelines (Thulke et al., 2024; Bulian et al., 2024), but sometimes resorts to comparisons (Mullappilly et al., 2023; Thulke et al., 2024) or simple user preferences (Mullappilly et al., 2023). Third, human evaluation can be influenced by factors other than the quality of the answer – for example, the observer-expectancy effect (Rosenthal, 1976). For instance, in ChatClimate (Vaghefi et al., 2023), humans evaluate the accuracy of text generated by a generic ChatGPT vs. the proposed ChatClimate chatbot. The text produced by the ChatClimate bot always gives specific references (down to page numbers), while the generic ChatGPT does not. This allows evaluators to easily identify the more precise answer based solely on the type of references used, raising the concern that this criterion may have disproportionately influenced assessments, potentially overlooking broader aspects such as overall accuracy or quality. LLM-based evaluation has its own challenges: Language models are known to reproduce human biases (Gallegos et al., 2024). They are sensitive to the way questions are formulated, to the order of answers in multiple-choice questions (Pezeshkpour and Hruschka, 2024), and to cues that can influence human evaluation such as using “Our Model”/“Their model” (which is the wording used in Mullappilly et al. (2023)). LLM-based evaluation should thus always adhere to the same rigorous standards that are applied to human evaluation.
Figure 1. Upset plot (Lex et al., 2014) of the methods used in the studies mentioned in this literature review. Each bar counts the number of studies reporting a given set of methods. The number of a papers using a given method is reported on the horizontal bars. We have highlighted in red the intersection sets that include both fine-tuned Transformers and keyword-based methods.
Baselines
Figure 1 shows the approaches that are commonly used in greenwashing detection tasks (more details in Table LABEL:tab:models). We see that the most common solution is fine-tuning transformer models; more recent approaches include using LLMs as few-shot learners. The figure also shows which baselines have been used:
  • Random/Majority Baseline: Any proposed approach should beat a random baseline. However, we find that only 9% of papers we studied compare their results to the random or majority baseline. For example, Friederich et al. (2021) reported performances under 50% for a binary classification task, which is worse than a theoretical random classifier on a balanced dataset. Their dataset is not balanced, but it is difficult to draw conclusions without the performance of a random classifier. Random classifiers cannot just suggest that a model underperforms, but also that a model has a comparatively good performance. For example, on SciDCC or ClimaTOPIC, the performance of models is around 50-60%. While this looks mediocre, it is in fact an excellent performance as the random classifier performed around 5-7% (Spokoyny et al., 2023). In case of a heavily imbalanced dataset, the majority baseline is more meaningful. If such a baseline is not provided, but the dataset is available, it can even be computed a posteriori. Unfortunately, neither is the case for the work of Friederich et al. (2021). Similarly, a constant value baseline can help understand the performance of a regression model. Sharma et al. (2024) and Bronzini et al. (2023) evaluated the prediction of the ESG score, but do not provide such a baseline. Thus, the reader is left to wonder whether a model that always predicts the mean value might not actually perform better.
  • Simple Baselines: If a keyword-based baseline performs well, or even better than a proposed method, then the proposed method is likely an overkill. For example, Rouen et al. (2023); Clarkson et al. (2020) showed that even for the prediction of an ESG score (which is a complex target), linguistic features such as the quantity and complexity of text are good predictors. This suggests that simple predictors could perform well even on complex tasks. Indeed, simple baselines can outperform finetuned approaches, for example on TCFD classification (Sampson et al., 2022) or on the ClimaINS dataset (Spokoyny et al., 2023). Beyond a validation (or invalidation) of the proposed approaches, a keyword-based baseline would also allow us to understand if models rely on an “understanding” of the text or merely on word cues. And yet, less than half of the works we studied report the performance of a keyword-based baseline.
  • Human Baselines: Using a human as a baseline can help determine whether a task is well-defined. If even a qualified human cannot achieve high accuracy, this most likely means that the task is subjective, ill-defined, or in disagreement with the dataset. Vice versa, if a human achieves high accuracy, but the model achieves even higher accuracy, this could mean that the model has acquired expert knowledge. However, with the exception of 3 studies, all studies discussed here not come with a human evaluation.
6.2. Model Robustness
When evaluating a model, it is important to know whether the performance on a dataset is due to an inherent capacity of the model or due to random factors that may be entirely different on another dataset.
Model Performance Evaluation
Many factors can influence the performance of a model (e.g. data sampling, hyper-parameters, random initialization, model architecture). Thus, a better performance in an experiment could be due also to these factors rather than due to the choice of the model. This is even more true when studies report differences in performance of less than 1%. For this reason, the uncertainty of the evaluation should always be quantified – by confidence intervals, p-values or at least the standard deviation. However, we found that only seven papers, out of the 61 we studied, quantify the uncertainty of the model performance (Webersinke et al., 2022; Leippold and Varini, 2020; Bingler et al., 2023; Schimanski et al., 2024a, 2023b; Lee and Kim, 2023; Schimanski et al., 2023a).
Noisy Text and Adversarial Attacks
Most studies mentioned in this survey trained and evaluated models on datasets that contain only clean examples. However, real-world data is usually noisy. Thus, the models may perform entirely differently on real-world datasets. Studies should quantify not just the performance of the models, but also their robustness, as fine-tuned models are prone to word-based adversarial attacks (Woloszyn et al., 2021).
Imbalanced Datasets
The class of interest is relatively rare in most tasks. For example, only 5% of text is in the positive class “climate-related” in the study of Yu et al. (2024). This means that (1) a simple baseline that always says “non-climate-related” will have a very high precision, and (2) a model trained on such a dataset would probably overfit. Therefore, most studies mentioned in this review trained and evaluated the models on relatively balanced datasets. On the other hand, a model evaluated only on a balanced dataset may fail when tested on real-world, unbalanced datasets (Friederich et al., 2021). For example, a model reaching an F1-score of 90% on a balanced dataset would significantly drop on an unbalanced dataset with 5% of text in the positive class (to a binary F1-score of 47%, and a macro F1-score of 71%). This is due to the large number of false positives. Thus, studies should consider both the balanced and the unbalanced case.
Precision-Recall Trade-off
In most methods (and most tasks), higher precision can be traded for smaller recall. In the case of greenwashing, both metrics can be useful: a higher recall is desirable when providing leads for human analysis, so as to avoid missing potential greenwashing cases. However, if the goal is to accurately measure greenwashing prevalence, minimizing false positives is critical, as false accusations can have serious repercussions. Most studies, however, are conducted without the use case in mind.
6.3. Data Sources
Close-source data and reproducibility
The studies mentioned in our review use various data sources, most of which are publicly available: annual and sustainable reports  (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023; Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022; Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023), social media (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024), specialized websites (Morio and Manning, 2023; Schimanski et al., 2023a; Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021). However, multiple works rely on closed-source data: ESG scores, carbon emissions, and machine-readable reports that are collected through private providers such as Refinitiv, Reuters, RavenPack, Trucost, Bloomberg, RobecoSAM, Compustat or Wharton (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022; Li et al., 2020; Clarkson et al., 2020; Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b; Vinella et al., 2023; Kraus et al., 2023). Relying on private providers limits the reproducibility of these works, for both academic work and applications.
No greenwashing datasets
The most fundamental shortcoming of current work on greenwashing is that there is no comprehensive real-world dataset with instances of greenwashing. This limits the conclusions that can be drawn from current studies drastically. The lack of a gold standard dataset can be explained by several factors:
  • The definitions of greenwashing are often vague and ambiguous, leaving room for interpretation, and making it difficult to annotate examples of greenwashing.
  • Green claims are often not self-contained, but spread across multiple paragraphs in a same document, across multiple documents, or even across different types of sources (such as tables, databases, etc.). Therefore, just collecting all the elements required to correctly identify greenwashing is a challenge on its own.
  • Understanding the claim and annotating greenwashing requires expertise, and sometimes even debate, hence producing greenwashing annotations is a difficult task.
  • Greenwashing is a strong accusation that might impact a company reputation, and that can also, if not substantiated by evidence, be considered a defamation.
To build a dataset of real-world greenwashing cases, several data collection strategies can be considered. To avoid relying on a subjective annotation, researchers can rely on a third-party judgment. For example, using reputable news articles that explicitly discuss greenwashing to find actual examples, drawing on prior research focused on detecting media mentions of greenwashing (Gourier and Mathurin, 2024). Another valuable resource are regulatory records from authorities such as the Advertising Standards Authority (ASA) or databases like greenwash.com.
Missing link to regulatory texts
Existing work often relies on its own definitions of greenwashing. This risks introducing subjectivity and incompatibility with other works. A safer choice is to use regulatory texts as a reference, as this is done, e.g., in Stammbach et al. (2023). Recently, the European Commission adopted the corporate sustainability reporting directive (CSRD) (Union, 2022), making environmental disclosure requirements explicit, and the directive on green claims (US Environmental Protection Agency, 2023), which defines requirements when making green claims. These could be good starting points for defining greenwashing in future research. Such regulations help not just with the definition of greenwashing, but also with its judgment. Instead of emitting a judgment, a method can link a statement to a regulatory text. For example, the works that predict TCFD simply link a statement to a recommendation (Ding et al., 2023; Auzepy et al., 2023; Brié et al., 2022; Bingler et al., 2022; Sampson et al., 2022). This approach makes the methods less prone to ad hoc definitions of scores and judgment strategies, and lets a human take the final decision.
Quantifying communication vs Quantifying information
With the exception of Bronzini et al. (2023), the studies reviewed in this work all measure how often a given text contains climate-related content, TCFD disclosures, or specific topics of interest. However, a company can repeatedly emphasize the same objective or repeatedly highlight a single initiative. With this, they skew the quantity of climate-related content in their favor, while they may still fail to implement any meaningful action. This phenomenon has been called “cheap talk” (Bingler et al., 2023). Future research should thus quantify not only how often some pieces of information are conveyed, but also how many different pieces there are.
7. Conclusion
In this survey, we have provided a comprehensive overview of current research initiatives on detection of greenwashing in textual corporate communications. Determining whether a communication contains greenwashing presents three fundamental challenges. First, the absence of a commonly accepted and actionable definition of greenwashing creates ambiguity, complicating any attempt at standardized assessment. Second, there are significant legal and reputational consequences of being accused of greenwashing. Finally, there are no explicit precedents that could serve as jurisprudence for systematic labeling. This makes the annotation of actual examples of greenwashing challenging, and indeed, there is currently no large-scale dataset of instances of greenwashing. For this reason, the detection of greenwashing has remained on a theoretical level without empirical validation (Moodaley and Telukdarie, 2023a). Most existing works target intermediate tasks such as identifying climate-relatedness or green claims. This work also serves its purposes: First, determining whether a text can contain greenwashing at all helps restricting attention to the relevant resources. Second, decomposing greenwashing into more manageable steps can lead to an overall better performance, as supported by the principles behind the Chain-of-Thought (Wei et al., 2022). Third, using frugal approaches helps reduce costs and energy consumption. Finally, intermediate tasks can help humans understand the final verdict, which is important for the explainability of decisions. Several challenges remain in these works:
  • Evaluation methodology. As models are capable of tackling more complex tasks, we need tools to correctly evaluate models on these complex tasks. This include building robust evaluation methods for generated texts or building classification benchmarks that are difficult to solve but simple to evaluate. Current results often lack hyper-parameter tuning, simple baselines, and inter-annotator agreement measures, making it difficult to draw definitive conclusions from them.
  • Model robustness. When evaluating models, we need to make sure to evaluate also their robustness in real-world applications that exhibit data quality issues or highly imbalanced classes. This includes quantifying the uncertainty of the performance measures. Current works often do not provide confidence intervals or robustness assessments.
  • Quantifying information. Most of the studies we reviewed quantify the amount of text related to a topic, rather than the amount of information. Future work should quantify also the information itself.
  • Relying on regulatory texts. Most existing works identify one or more supposed characteristics of a misleading claim, but do not confront the statements to regulatory texts. As regulatory texts are becoming more precise about the requirements of environmental communications, they will increasingly become the gold standard for identifying problematic claims.
Any step toward understanding the characteristics and formalizing greenwashing is a step toward developing AI systems capable of automatically identifying these misleading communications. While previous research has laid a foundation for the theoretical understanding of greenwashing, and for the detection of signals that can indicate misleading climate-related communications, future research should focus on building a representative and diverse corpus of real-world cases of potential greenwashing to first evaluate models and, secondly, to empirically analyze its mechanisms and patterns. We hope that future research will help identify and combat greenwashing more reliably, thereby fostering transparency and accountability in climate communication.
Limitations
This survey provides a comprehensive overview of studies on automatic detection of misleading climate-related corporate communications and its related tasks as of early 2025. A number of limitations remain: Given the extensive scope of the field, our review concentrated on NLP-based research, providing only limited discussion of topics like ESG prediction, which are predominantly explored through quantitative methods. Furthermore, our review relied primarily on keyword searches on arXiv, Google Scholar, and ResearchGate, complemented by references from identified papers. In this, we may have overlooked relevant studies that use alternative terminology or are published in less accessible venues. Finally, much of greenwashing happens not on the text level, but on the factual level (with claims that contradict data) or on the multi-modal level (across images or videos). These remain out of scope for our survey.
Acknowledgment
We would like to express our sincere gratitude to Nina Godart and AEF Info for their invaluable contributions and knowledge on the topic. We also extend our thanks to Takaya Sekine CFA for his insightful comments. This work was partially supported by the grant ANR-20-CHIA-0012-01 (“NoRDF”).
Appendix
Appendix A TCFD recommendations
The recommendations are directly taken from the TCFD website: https://www.fsb-tcfd.org/recommendations/
Governance
Disclose the organization’s governance around climate-related risks and opportunities.
  • Describe the board’s oversight of climate-related risks and opportunities.
  • Describe management’s role in assessing and managing climate-related risks and opportunities.
Strategy
Disclose the actual and potential impacts of climate-related risks and opportunities on the organization’s businesses, strategy, and financial planning where such information is material.
  • Describe the climate-related risks and opportunities the organization has identified over the short, medium, and long term.
  • Describe the impact of climate-related risks and opportunities on the organization’s businesses, strategy, and financial planning.
  • Describe the resilience of the organization’s strategy, taking into consideration different climate-related scenarios, including a 2°C or lower scenario.
Risk Management
Disclose how the organization identifies, assesses, and manages climate-related risks.
  • Describe the organization’s processes for identifying and assessing climate-related risks.
  • Describe the organization’s processes for managing climate-related risks.
  • Describe how processes for identifying, assessing, and managing climate-related risks are integrated into the organization’s overall risk management.
Metrics and Targets
Disclose the metrics and targets used to assess and manage relevant climate-related risks and opportunities where such information is material.
  • Disclose the metrics used by the organization to assess climate-related risks and opportunities in line with its strategy and risk management process.
  • Disclose Scope 1, Scope 2 and, if appropriate, Scope 3 greenhouse gas (GHG) emissions and the related risks.
  • Describe the targets used by the organization to manage climate-related risks and opportunities and performance against targets.
A.1. Data sources
Source Papers
10-K reports (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023)
Manually collected reports (Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022; Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023)
Refinitiv (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022)
Reuters (Li et al., 2020; Clarkson et al., 2020)
ESG score private providers (Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b)
CDP (Spokoyny et al., 2023)
Open for Good Data (Vinella et al., 2023)
ClimateWatch (Kraus et al., 2023)
Twitter/Reddit (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024)
LobbyMap (Morio and Manning, 2023)
Net-zero tracker (Schimanski et al., 2023a)
Fact-checking websites (Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021)
Table 10. Data Sources
Publicly available Reports
Many studies rely on 10-K reports (Doran and Quinn, 2009; Kölbel et al., 2021; Chou et al., 2023; Huang et al., 2020; Kraus et al., 2023) as they are easily accessible (EDGAR), uniformly structured, and mandatory for companies listed on US financial markets. Other reports, such as sustainable reports, are not centralized and mandatory, therefore they need to be collected from other sources. They are often collected manually from the company’s website (Kheradmand et al., 2021; Rouen et al., 2023; Lai and Chen, 2023; Kang and Kim, 2022; Auzepy et al., 2023; Brié et al., 2022; Webersinke et al., 2022), or from other websites (Luccioni et al., 2020; Kheradmand et al., 2021; Bronzini et al., 2023; Lee and Kim, 2023; Polignano et al., 2022; Kraus et al., 2023) such as from the GRI, SASB, TCFD, ResponsibilityReports, AnnualReports, Korean Standards Association. Polignano et al. (2022) relied on Sustainability Disclosure Database from GRI. To the best of our knowledge, it is not available any longer.
The reports are openly available, however they are most of the time in a PDF format, and must be parsed to produce machine-readable versions. While Khan (2024) proposed a machine-readable dataset for the 10-K reports, it is often not the case for other types of reports. (Bronzini et al., 2023; Polignano et al., 2022; Kang and Kim, 2022) provide a sample of processed reports.
Private Providers
Machine-readable reports, as well as earning call transcripts are often provided through a private data provider such as Refinitiv (Sautner et al., 2023; Bingler et al., 2023; Stammbach et al., 2023; Friederich et al., 2021; Brié et al., 2022; Lee and Raschke, 2023; Webersinke et al., 2022) or Reuters (Li et al., 2020; Clarkson et al., 2020). Similarly, financial data (e.g. ESG score) is often collected through data providers such as RavenPack or S&P Global Trucost or Wharton’s research platform WRDS, Reuters, Bloomberg, Refinitiv, RobecoSAM and Compustat (Li et al., 2020; Sautner et al., 2023; Mehra et al., 2022; Sharma et al., 2024; Clarkson et al., 2020; Schimanski et al., 2023b)
Carbon emission data
Carbon emissions of companies can be collected through the private data providers, or through CDP (Spokoyny et al., 2023), Open for Good Data (Vinella et al., 2023) or ClimateWatch website (Kraus et al., 2023).
Media and Social Media
As companies communicate with the public mainly through marketing and social media, other data sources include Twitter and Reddit (Woloszyn et al., 2021; Luo et al., 2020; Vaid et al., 2022; Coan et al., 2021; Oppong-Tawiah and Webster, 2023; Sharma et al., 2024).
Climate expert websites
Finally, projects such as LobbyMap or Net Zero Tracker, that track corporate policy and engagements, can help build annotated datasets (Morio and Manning, 2023; Schimanski et al., 2023a). Similarly, relying on a fact-checking website such as https://climatefeedback.org/ or www.skepticalscience.com is an efficient way of building a dataset (Jin et al., 2022; Leippold and Varini, 2020; Coan et al., 2021)
Appendix B Detailed model list
Table 11. Solutions proposed in each paper
Paper Dataset Baselines BERT Sent. Sim. Other
Pretraining Models on Climate-Related Text
Webersinke et al. (2022) Domain-specific corpus - Pre-Training ClimateBERT
Vaghefi et al. (2022) Deep Climate Change - Pre-Training ClimateGPT-2
Mullappilly et al. (2023) Deep Climate Change - Pre-Training stsb-xlm-r-multilingual Arabic Mini-ClimateGPT, Vicuna, Alpaca, Dolly V2, Alpaca ar
Schimanski et al. (2023b) Domain-specific corpus - Pre-Training EnvironmentalBERT
Yu et al. (2024) climateBUG-data - Pre-Training climateBUG-LM
Thulke et al. (2024) Domain-specific corpus - Pre-Training ClimateGPT-7/13/70B
Thulke et al. (2024) ClimaBENCH - Evaluation Stability-3B, Pythia-6.9B, Falcon-7B, Mistral-7B, Llama-2-7B, Jais-13B, Jais-13B-Chat, Llama-2-7/13/70B, Llama-2-Chat-7/13/70B, ClimateGPT-7/13/70B
Climate-related topic detection
Doran and Quinn (2009) 10-K filings - climate Keywords
Engle et al. (2020) WSJ - climate TF-IDF cosine similarity
Leippold and Varini (2020) climatext NB, Keywords BERT
Garrido-Merchán (Garrido-Merchán et al., 2023) climatext ClimateBERT, BERT
Spokoyny et al. (2023) climatext Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Webersinke et al. (2022) climate detection distilRoBERTa, climateBERT
Bingler et al. (2023) climate detection NB, SVM + BoW, SVM + ELMo climateBERT
Brié et al. (2022) climate detection Random Forest climateBERT, FinBERT
Yu et al. (2024) climateBUG data BERT, FinBERT, ClimateBERT, climateBUG-LM
Lin et al. (2023b) sustainable signals review
Climate-related subtopic detection
Kang and Kim (2022) Sustainable Reports - topics sentence-BERT
Mishra and Mittal (2021) sciDCC
Spokoyny et al. (2023) sciDCC Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Vaid et al. (2022) climateEng BERT, RoBERTa, DistilBERT FastText
Spokoyny et al. (2023) climateEng Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Jain et al. (2023) ledger - topics TF-IDF BERT, ClimateBERT and RoBERTa. sentence-BERT Word2Vec-based Random Forest
Schimanski et al. (2024a) ESGBERT-nature Keywords ClimateBERT, RoBERTa, DistilRoBERTa, EnvironmentalBERT
Spokoyny et al. (2023) ClimaTOPIC Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Detecting climate-related financial disclosure
Ding et al. (2023) Reports - Disclosure TF-IDF cosine similarity
Auzepy et al. (2023) Reports - TCFD BART+MNLI
Brié et al. (2022) Sustainable Reports structural topic model (STM)
Bingler et al. (2022) climate tcfd rec. TF-IDF + random forest RoBERTa, ClimateBERT (RoBERTa+LogReg) Sentence Enc.
Sampson et al. (2022) Reports - TCFD TF-IDF ClimateBERT, distilRoBERTa kNN (sentence-transformers embeddings), stacked models
In-depth Disclosure: Climate Risk Classification
Li et al. (2020) Reports - Risk dictionaries (risk)
Chou et al. (2023) 10-K filings - Risk 3 dictionaries (risk: Climate, Impact, Litigious)
Kheradmand et al. (2021) Annual/Sustainability Reports - Risk dictionaries (risk)
Sautner et al. (2023) earning calls - Risk dictionaries (risk/opportunity)
Polignano et al. (2022) Sustainability Reports - Sentiment TextBlob, Sentita
Kang and Kim (2022) Sustainability Reports - Sentiment DistilBERT (sentiment analysis)
Kölbel et al. (2021) 10-K filings - Risk BoW, TF-IDF BERT
Friederich et al. (2021) Reports - Risk SVM distilBERT, RoBERTa
Bingler et al. (2023) climate sentiment NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Webersinke et al. (2022) climate sentiment distilRoBERTa, climateBERT
Xiang and Fujii (2023) Web of Science, Scopus - Sentiment BERT, Domain-specific BERT, TinyBERT, DistilBERT, ClimateBERT CCLA+Max-Pooling, Bi-LSTM-Attention, POS-Bi-LSTM-Attention, DARE
Detecting environmental, social, and governance disclosure
Rouen et al. (2023) Sustainability Reports - ESG industry-topic dictionaries
Mehra et al. (2022) Sustainability Reports - ESG sentence BERT, Universal Sentence Encoder, and Deep Averaging Network
Huang et al. (2020) Sustainability Report, 10-K filings - ESG NB, SVM, RF FinBERT, BERT CNN, LSTM
Schimanski et al. (2023b) ESGBERT E,S,G BERT, RoBERTa, DistilRoBERTa, E/S/G-RoBERTa, E/S/G-DistilRoBERTa
Lee and Kim (2023) Korean Reports - ESG Multilingual BERT, KoBERT, DistilKoBERT, KB-ALBERT, KLUE-RoBERTa
Bronzini et al. (2023) Sustainability Reports - ESG INSTRUCTOR-xl
Green claim detection
Woloszyn et al. (2021) green claims RoBERTa, BERTweet Flair
Stammbach et al. (2023) environmental claims Random, Majority, TF-IDF + SVM, Character n-gram SVM DistilBERT, ClimateBERT, RoBERTa ClaimBuster RoBERTa, Pledge Detection RoBERTa
Lin et al. (2023b) Product Description - Claim environmental claim detection (finetuned distilBERT)
Green claim characteristics
Bingler et al. (2023) climate-commitments-actions NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Schimanski et al. (2023a) netzero/reduction ClimateBERT, DistilRoBERTa, RoBERTa GPT-3.5-turbo
Woloszyn et al. (2021) green claims 3 RoBERTa, BERTweet Flair
Bingler et al. (2023) climate specificity NB, SVM + BoW, SVM + ELMo distilRoBERTa, climateBERT
Green stance detection
Morio (Morio and Manning, 2023) LobbyMap Majority, TF-IDF BERT, RoBERTa, ClimateBERT, Longformer
Luo et al. (2020) DeSMOG
Vaid et al. (2022) ClimateStance BERT, RoBERTa, DistilBERT FastText
Lai and Chen (2023) Reports - Stance BERT, RoBERTa, ClimateBERT
Spokoyny et al. (2023) ClimateStance Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Spokoyny et al. (2023) DeSMOG Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Leippold (Leippold and Diggelmann, 2020) ClimateFEVER ALBERT(FEVER)
Wang et al. (2021) ClimateFEVER RoBERTa (ClimateFEVER) RoBERTa (FEVER), RoBERTa (ClimateFEVER) + UDA
Spokoyny et al. (2023) ClimateFEVER Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Vaghefi et al. (2022) ClimateFEVER GPT-2, climateGPT-2
Xiang and Fujii (2023) ClimateFEVER BERT, RoBERTa, DistilRoBERTa, ClimateBERT, POS-Bi-LSTM-Attention
Webersinke et al. (2022) ClimateFEVER distilRoBERTa, climateBERT
Vaid et al. (2022) ClimateFEVER BERT, RoBERTa, DistilBERT FastText
Question Answering
Luccioni et al. (2020) Reports - QA RoBERTa
Spokoyny et al. (2023) ClimaQA BM25 MiniLM
Spokoyny et al. (2023) ClimaINS Random, Majority, SVM BERT, RoBERTa, DistilRoBERTa, Longformer, SciBERT, ClimateBERT
Rony et al. (2022) CCMRC - QA pairs ALBERT (trained on SQuAD then finetuned) Sentence-BERT
Schimanski et al. (2023a) Net-Zero Tracker - Target RoBERTa Quad2
Ni et al. (2023) Reports - RAG ChatGPT
Vaghefi et al. (2023) Reports - RAG - Human Eval GPT-4
Mullappilly et al. (2023) Clima500-Instruct - QA instructions Vicuna (Mini-ClimateGPT)
Thulke et al. (2024) Instruction following corpus - Human Eval ClimateGPT-70B, ClimateGPT-7B, ClimateGPT-FSC-7B
Thulke et al. (2024) ClimRetrieve ColBERTv2, DRAGON, GTE-base, text-embedding-ada-002, text-embedding-3-small, text-embedding-3-large BM25
Deceptive Techniques
Oppong (Oppong-Tawiah and Webster, 2023) Tweets - Fake News linguistic indicators
Coan et al. (2021) contrarian claims Logistic, SVM RoBERTa ULMFiT (over sample, weighted, focal Loss, logisitic, SVM), RoBERTa + Logistic
Jin et al. (2022) logicClimate ELECTRA, BERT lectra-StructureAware
Bhatia et al. (2021) climate change scepticism - Neutralization SVM BERT MTEXT
Environmental performance prediction
Jain et al. (2023) Ledger - Scope 3 Emissions Alpaca, WizardLM, Flan-T5, ChatGPT
Bronzini et al. (2023) Sustainable Reports - ESG OLS
Mehra et al. (2022) Reports 10-Qs - ESG Majority BERT, ESGBERT
Clarkson et al. (2020) CSR Reports - CSR performer linguistic features + random forest, XGBoost, SVM
Lin et al. (2023b) Product Description / Review - Finch Score Lasso, Gradient Boosting NOCATE + transformer, SUSTAINABLESIGNALS + transformer
Sharma et al. (2024) Multimodel Tweet - ESG CLIP+MLP, Sentence-BERT+MLP
Appendix C Reported Performances
We have reported all the performances computed in each study when available. We did not include the performance table if the original study did not perform an evaluation (for example, in zero-short/unsupervised settings) or if the evaluation is a heavily modified metric (for example, when using a quantitative measure alongside other signals to compute a financial factor). We only reported traditional NLP metrics. When available, we reported the F1-score, however, if the F1-score is not computed in the study, we reported the available metrics. The computation methods of the F1-score (macro, micro, weighted) are specified when enough information is given to state it confidently. We reported one table per dataset, which might contain multiple tasks, but multiple studies can tackle the same dataset and are then grouped to be able to compare the results.
C.1. Climate-related Topic Classification
Source Model F1
Webersinke et al. (2022) distilRoBERTa 98.6% ±$\pm 1$
Webersinke et al. (2022) ClimateBERT 99.1% ±$\pm 0.8$
Bingler et al. (2023) Naive Bayes 87%
Bingler et al. (2023) SVM + BoW 87%
Bingler et al. (2023) SVM + ELMo 89%
Bingler et al. (2023) ClimateBERT 97%
Brié et al. (2022) ClimateBERT 98.59%
Brié et al. (2022) FinBERT 96.67%
Brié et al. (2022) Random Forest 91.65%
(a) F1-scores (weighted average) on ClimateBERT’s climate detection dataset (Bingler et al., 2023). For Webersinke et al. (2022) scores are displayed as mean ±$\pm std$. Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
Source Model F1
Yu et al. (2024) BERT 90.81%
Yu et al. (2024) FinBERT 90.82%
Yu et al. (2024) ClimateBERT 91.07%
Yu et al. (2024) climateBUG-LM 91.36%
(b) F1-scores (macro average) on ClimateBUG-data (Yu et al., 2024).
Source Model Wikipedia 10-K Claims
F1 (bin.) F1 (bin.) F1 (bin.)
Leippold and Varini (2020) best BERT 80% ±$\pm 6$ 95% $\pm$ 2 83% $\pm$ 1
Leippold and Varini (2020) best NB 60% ±$\pm 7$ 90% ±$\pm 3$ 72% ±$\pm 2$
Leippold and Varini (2020) best Keywords 67% ±$\pm 7$ 88% ±$\pm 3$ 70% ±$\pm 1$
Leippold and Varini (2020) BERT-AL-wiki 69% ±$\pm 6$ 71% ±$\pm 4$ 81% ±$\pm 1$
Garrido-Merchán (Garrido-Merchán et al., 2023) BERT-AL-wiki - 91% ±$\pm 0.6$
Garrido-Merchán (Garrido-Merchán et al., 2023) climateBERT-AL-wiki - 93% ±$\pm 0.7$
Combined - F1 (macro)
Spokoyny et al. (2023) Majority 42.08%
Spokoyny et al. (2023) Random 46.86%
Spokoyny et al. (2023) SVM 83.39%
Spokoyny et al. (2023) BERT 87.04%
Spokoyny et al. (2023) RoBERTa 85.97%
Spokoyny et al. (2023) DistilRoBERTa 86.06%
Spokoyny et al. (2023) Longformer 87.80%
Spokoyny et al. (2023) SciBERT 83.29%
Spokoyny et al. (2023) ClimateBERT 85.14%
(c) F1-scores on ClimaText (Leippold and Varini, 2020). For Leippold and Varini (2020) scores are displayed as mean ±$\pm std$.
Table 12. Reported performances for the datasets on climate-related topic detection (which include climate, sustainability and environment).
C.2. Thematic Analysis
Source Model Prec.
Bingler et al. (2022) TF-IDF+random forest 24%
Bingler et al. (2022) Sentence Enc. 23%
Bingler et al. (2022) RoBERTa Para. 22%
Bingler et al. (2022) RoBERTa Sent. 75%
Bingler et al. (2022) ClimateBERT* 84%
Bingler et al. (2022) ClimateBERT* blahblF1: 84%
(a) Average Precisions and F1-score (macro average) on TCFD recommendations classification (Bingler et al., 2022). (*They refer to a finetuned RoBERTa with a Logistic Regression as ClimateBERT)
Source Model ROC-AUC
Sampson et al. (2022) TF-IDF (1) 0.867
Sampson et al. (2022) kNN (2) 0.847
Sampson et al. (2022) ClimateBERT (3) 0.852
Sampson et al. (2022) distilRoBERTa (4) 0.819
Sampson et al. (2022) (1) + (2) 0.884
Sampson et al. (2022) (1) + (3) 0.865
Sampson et al. (2022) (1) + (4) 0.857
(b) ROC-AUC scores on TCFD classification task (Sampson et al., 2022)
Source Model F1
Auzepy et al. (2023) BART+MNLI 56.68%
(c) F1-scores (macro average) on TCFD classification (Auzepy et al., 2023)
Table 13. Reported performances for the datasets on TFCD classification.
Source Model F1
Huang et al. (2020) FinBERT 89.6%
Huang et al. (2020) BERT 86.9%
Huang et al. (2020) NB 75.0%
Huang et al. (2020) SVM 75.0%
Huang et al. (2020) RF 78.3%
Huang et al. (2020) CNN 81.9%
Huang et al. (2020) LSTM 86.0%
(a) F1-scores (micro average) on ESG classification (Huang et al., 2020)
Source Model Accuracy
Lee and Kim (2023) Multiling. BERT 84.09% ±$\pm 0.092$
Lee and Kim (2023) KoBERT 85.98% ±$\pm 0.163$
Lee and Kim (2023) DistilKoBERT 84.89% ±$\pm 0.081$
Lee and Kim (2023) KB-ALBERT 85.62% ±$\pm 0.026$
Lee and Kim (2023) KLUE (base) 86.05% ±$\pm 0.125$
Lee and Kim (2023) KLUE (large) 86.66% ±$\pm 0.099$
(b) Accuracy on ESG classification (Lee and Kim, 2023). Accuracy reported as mean ±$\pm std$
Environment F1
Schimanski et al. (2023b) BERT 91.85% ±$\pm 1.25$
Schimanski et al. (2023b) RoBERTa 92.35% ±$\pm 2.29$
Schimanski et al. (2023b) DistilRoBERTa 90.97% ±$\pm 2.00$
Schimanski et al. (2023b) EnvRoBERTa 93.19% ±$\pm 1.40$
Schimanski et al. (2023b) EnvDistilRoBERTa 92.35% ±$\pm 1.65$
Social
Schimanski et al. (2023b) BERT 89.84% ±$\pm 0.79$
Schimanski et al. (2023b) RoBERTa 89.87% ±$\pm 1.35$
Schimanski et al. (2023b) DistilRoBERTa 90.59% ±$\pm 1.03$
Schimanski et al. (2023b) SocRoBERTa 91.90% ±$\pm 1.79$
Schimanski et al. (2023b) SocDistilRoBERTa 91.24% ±$\pm 1.86$
Governance
Schimanski et al. (2023b) BERT 76.86% ±$\pm 1.77$
Schimanski et al. (2023b) RoBERTa 77.03% ±$\pm 1.82$
Schimanski et al. (2023b) DistilRoBERTa 76.65% ±$\pm 2.39$
Schimanski et al. (2023b) GovRoBERTa 78.48% ±$\pm 2.62$
Schimanski et al. (2023b) GovDistilRoBERTa 78.86% ±$\pm 1.59$
(c) F1-scores (unspecified) on environment, social, and governance datasets (Schimanski et al., 2024a). F1-scores are reported as mean ±$\pm std$
Table 14. Reported performances for the datasets on ESG classification
Source Model Water Forest Biodiv. Nature
Schimanski et al. (2024a) EnvironmentalBERT 94.47% ±$\pm 1.37$ 95.37% ±$\pm 0.92$ 92.76% ±$\pm 1.01$ 94.19% ±$\pm 0.81$
Schimanski et al. (2024a) ClimateBERT 95.10% ±$\pm 1.13$ 95.34% ±$\pm 0.94$ 92.49% ±$\pm 1.03$ 93.50% ±$\pm 0.64$
Schimanski et al. (2024a) RoBERTa 94.55% ±$\pm 0.86$ 94.78% ±$\pm 0.48$ 92.46% ±$\pm 1.54$ 93.97% ±$\pm 0.26$
Schimanski et al. (2024a) DistilRoBERTa 94.98% ±$\pm 1.16$ 95.29% ±$\pm 0.65$ 92.29% ±$\pm 1.23$ 93.55% ±$\pm 0.72$
Schimanski et al. (2024a) Keywords - - 63.03% 61.00%
(a) F1-scores (unspecified) on Nature-related topic classification (Schimanski et al., 2024a). F1-scores are reported as mean ±$\pm std$.
Source Model F1
Spokoyny et al. (2023) Majority 0.79%
Spokoyny et al. (2023) Random 5.05%
Spokoyny et al. (2023) SVM 48.02%
Spokoyny et al. (2023) BERT 54.74%
Spokoyny et al. (2023) RoBERTa 52.90%
Spokoyny et al. (2023) DistilRoBERTa 51.13%
Spokoyny et al. (2023) Longformer 54.79%
Spokoyny et al. (2023) SciBERT 51.83%
Spokoyny et al. (2023) ClimateBERT 52.97%
(b) F1-scores (macro average) on SciDCC (Mishra and Mittal, 2021).
Source Models F1
Spokoyny et al. (2023) Majority 3.65%
Spokoyny et al. (2023) Random 6.45%
Spokoyny et al. (2023) SVM 58.34%
Spokoyny et al. (2023) BERT 64.64%
Spokoyny et al. (2023) RoBERTa 65.22%
Spokoyny et al. (2023) DistilRoBERTa 63.61%
Spokoyny et al. (2023) Longformer 64.03%
Spokoyny et al. (2023) SciBERT 63.62%
Spokoyny et al. (2023) ClimateBERT 64.24%
(c) F1-scores (macro average) on the climaTOPIC (Spokoyny et al., 2023).
Source Model F1
Jain et al. (2023) all-mpnet-base-v2 43.7%
Jain et al. (2023) TF-IDF 69%
Jain et al. (2023) Word2Vec 72%
Jain et al. (2023) RoBERTa 87.19%
Jain et al. (2023) BERT 87.14%
Jain et al. (2023) ClimateBERT 85.20%
(d) F1-scores (weighted average) on ledger transaction classification (Jain et al., 2023).
Source Model F1
Vaid et al. (2022) FastText 63.8%
Vaid et al. (2022) BERT-Base 69.6%
Vaid et al. (2022) BERT-Large 69.5%
Vaid et al. (2022) RoBERTa-Base 73.4%
Vaid et al. (2022) RoBERTa-Large 73.5%
Vaid et al. (2022) DistilBERT 69.4%
Spokoyny et al. (2023) Majority 13.83%
Spokoyny et al. (2023) Random 16.71%
Spokoyny et al. (2023) SVM 51.81%
Spokoyny et al. (2023) BERT 71.78%
Spokoyny et al. (2023) RoBERTa 74.58%
Spokoyny et al. (2023) DistilRoBERTa 72.33%
Spokoyny et al. (2023) Longformer 72.28%
Spokoyny et al. (2023) SciBERT 70.50%
Spokoyny et al. (2023) ClimateBERT 71.83%
(e) F1-scores (macro average) on ClimateEng (Vaid et al., 2022).
Table 15. Reported performances for the datasets on climate-related topic detection (other topics.)
C.3. In-depth Disclosure: Climate Risk Classification
Binary 2 Classes 5 Classes
Source Model REALISTIC REALISTIC REALISTIC HARD NEG. DISCRIMINATORY
Friederich et al. (2021) SVM 29.0% 35.1% 20.4% 45.7% 59.9%
Friederich et al. (2021) distilBERT 44.4% 49.7% 24.1% 43.1% 55.8%
Friederich et al. (2021) RoBERTa 49.6% 44.6% 35.6% 52.8% 59.6%
(a) F1-scores (macro average) for climate risk disclosure Friederich et al. (2021).
Source Model F1
Xiang and Fujii (2023) BERT-Base 93.1%
Xiang and Fujii (2023) BERT-Base (Domain-spe.) 95.5%
Xiang and Fujii (2023) TinyBERT 87.0%
Xiang and Fujii (2023) DistilBERT 89.9%
Xiang and Fujii (2023) ClimateBERT 87.5%
Xiang and Fujii (2023) CCLA+Max-Pooling 82.9%
Xiang and Fujii (2023) Bi-LSTM-Attention 71.9%
Xiang and Fujii (2023) POS-Bi-LSTM-Attention 88.2%
Xiang and Fujii (2023) DARE 89.4%
(b) F1-scores (macro average) on Risk/Opportunity classification (Xiang and Fujii, 2023).
Source Model F1
Webersinke et al. (2022) distilRoBERTa 82.5% $\pm$ 4.6
Webersinke et al. (2022) climateBERT 83.8% $\pm$ 3.6
Bingler et al. (2023) Naive Bayes 72%
Bingler et al. (2023) SVM + BoW 72%
Bingler et al. (2023) SVM + ELMo 75%
Bingler et al. (2023) climateBERT 80%
(c) F1-scores (weighted average) on risk/opportunity detection Webersinke et al. (2022); Bingler et al. (2023). Webersinke et al. (2022)’s F1-scores are reported as mean ±$\pm std$. Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not shown in this table for simplicity.
Source Model F1
Kölbel et al. (2021) Bag-of-Words 83.63%
Kölbel et al. (2021) TF-IDF 80.05%
Kölbel et al. (2021) BERT 94.66%
(d) F1-scores (unspecified) for risk classification (General/Transition risk/Physical Risk) (Kölbel et al., 2021).
Table 16. Reported performances for the datasets on risk detection and classification.
C.4. Green Claim Detection
Source Model F1
Stammbach et al. (2023) Majority baseline 00.0%
Stammbach et al. (2023) Random baseline 33.5%
Stammbach et al. (2023) ClaimBuster RoBERTa 33.9%
Stammbach et al. (2023) Pledge Detection RoBERTa 26.4%
Stammbach et al. (2023) TF-IDF SVM 69.1%
Stammbach et al. (2023) Character n-gram SVM 70.9%
Stammbach et al. (2023) DistilBERT 83.7%
Stammbach et al. (2023) ClimateBERT 83.8%
Stammbach et al. (2023) RoBERTa_base 82.4%
Stammbach et al. (2023) RoBERTa_large 84.9%
(a) F1-scores (binary) on environmental claim detection (Stammbach et al., 2023).
Source Model F1
Woloszyn et al. (2021) Bin RoBERTa Bal 92.08%
Woloszyn et al. (2021) Bin BERTweet Bal 91.48%
Woloszyn et al. (2021) Bin Flair Bal 88.52%
Woloszyn et al. (2021) Bin RoBERTa Unbal 88.88%
Woloszyn et al. (2021) Bin BERTweet Unbal 84.30%
Woloszyn et al. (2021) Bin Flair Unbal 81.55%
(b) F1-scores (binary) on green claim detection (Woloszyn et al., 2021).
Table 17. Reported performances for the datasets on green claim detection.
C.5. Green Claim Characteristics
Source Model F1
Woloszyn et al. (2021) Mult RoBERTa Bal. 81.45%
Woloszyn et al. (2021) Mult BERTweet Bal. 75.14%
Woloszyn et al. (2021) Mult Flair Bal. 69.96%
Woloszyn et al. (2021) Mult RoBERTa Unbal. 76.82%
Woloszyn et al. (2021) Mult BERTweet Unbal. 59.71%
Woloszyn et al. (2021) Mult Flair Unbal. 38.15%
(a) F1-scores (unspecified) on Implicit/Explicit green claim detection (Woloszyn et al., 2021).
Source Model Accuracy
Schimanski et al. (2023a) ClimateBERT 96.2% ±$\pm 0.4$
Schimanski et al. (2023a) DistilRoBERTa 94.4% ±$\pm 0.7$
Schimanski et al. (2023a) RoBERTa-base 95.8% ±$\pm 0.6$
Schimanski et al. (2023a) GPT-3.5-turbo 92.0%
(b) Accuracy on net-zero/reduction classification (Schimanski et al., 2023a). Accuracy is reported as mean ±$\pm std$.
Source Model Commit. & Act. Specificity
Bingler et al. (2023) Naive Bayes 75% 75%
Bingler et al. (2023) SVM + BoW 76% 75%
Bingler et al. (2023) SVM + ELMo 79% 76%
Bingler et al. (2023) climateBERT 81% 77%
(c) F1-scores (weighted average) on Commitment&Action and Specificity classification (Bingler et al., 2023). Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
Table 18. Reported performances for the datasets on claim characteristics classification.
C.6. Green Stance Detection
Source Model F1
Vaghefi et al. (2022) GPT-2 67%
Vaghefi et al. (2022) climateGPT-2 72%
Claim
Leippold and Diggelmann (2020) ALBERT(FEVER) 32.85%
Xiang and Fujii (2023) BERT 80.7%
Xiang and Fujii (2023) RoBERTa 72.3%
Xiang and Fujii (2023) DistilRoBERTa 72.0%
Xiang and Fujii (2023) ClimateBERT 76.8%
Xiang and Fujii (2023) POS-Bi-LSTM-Att. 79.3%
Webersinke et al. (2022) DistilRoBERTa 74.8% ±$\pm 3.6$
Webersinke et al. (2022) ClimateBERT 75.7% ±$\pm 4.4$
Subset 95 examples
Wang et al. (2021) RoBERTa (FEVER) 49.2%
Wang et al. (2021) + ClimateFEVER 64.2%
Wang et al. (2021) + UDA ratio 18 71.8%
Evidence
Spokoyny et al. (2023) Majority 26.08%
Spokoyny et al. (2023) Random 30.62%
Spokoyny et al. (2023) BERT 62.47%
Spokoyny et al. (2023) RoBERTa 60.74%
Spokoyny et al. (2023) DistilRoBERTa 61.54%
Spokoyny et al. (2023) Longformer 60.82%
Spokoyny et al. (2023) SciBERT 62.68%
Spokoyny et al. (2023) ClimateBERT 61.54%
(a) F1-score (macro average) on ClimateFEVER Leippold and Diggelmann (2020). The prediction can be made at Claim level or for each Claim-Evidence pair (Evidence). Webersinke et al. (2022)’s F1-scores are reported as mean ±$\pm std$.
Source Model Acc. F1
Luo et al. (2020) Majority class 0.43 17%
Luo et al. (2020) Linear 0.62 60%
Luo et al. (2020) BERT 0.75 73%
Luo et al. (2020) Human 0.71 -
(b) Accuracy and F1-scores (macro average) on GWSD (Luo et al., 2020).
Source Model F1
Vaid et al. (2022) FastText 34.3%
Vaid et al. (2022) BERT-Base 46.4%
Vaid et al. (2022) BERT-Large 48.9%
Vaid et al. (2022) RoBERTa-Base 51.0%
Vaid et al. (2022) RoBERTa-Large 48.9%
Vaid et al. (2022) DistilBERT 44.8%
Spokoyny et al. (2023) Majority 29.68%
Spokoyny et al. (2023) Random 25.52%
Spokoyny et al. (2023) SVM 42.92%
Spokoyny et al. (2023) BERT 55.37%
Spokoyny et al. (2023) RoBERTa 59.69%
Spokoyny et al. (2023) DistilRoBERTa 52.51%
Spokoyny et al. (2023) Longformer 34.68%
Spokoyny et al. (2023) SciBERT 48.67%
Spokoyny et al. (2023) ClimateBERT 52.84%
(c) F1-scores (macro average) on ClimateStance (Vaid et al., 2022).
Source Model macro bin.
Baselines
Lai and Chen (2023) BERT - 58%
Lai and Chen (2023) RoBERTa - 89%
Lai and Chen (2023) ClimateBERT 90% 90%
SeqGAN
Lai and Chen (2023) BERT - 83%
Lai and Chen (2023) RoBERTa - 87%
Lai and Chen (2023) ClimateBERT 90% 90%
MaliGAN
Lai and Chen (2023) BERT - 74%
Lai and Chen (2023) RoBERTa - 88%
Lai and Chen (2023) ClimateBERT 91% 91%
RankGAN
Lai and Chen (2023) BERT - 82%
Lai and Chen (2023) RoBERTa - 87%
Lai and Chen (2023) ClimateBERT 91% 90%
(d) F1-scores (binary and macro average) on stance on climate change remediation efforts) (Lai and Chen, 2023).
Table 19. Reported performance for the datasets on stance detection (Part 1)
Document Page overlap Strict
Source Model P Q S P Q S P Q S
Morio and Manning (2023) Most-frequent 46.7% 52.6% 36.8% 51.8% 25.6% 19.8% 41.2% 19.6% 17.5%
Morio and Manning (2023) Linear 66.0% 61.9% 50.3% 71.4% 44.5% 36.1% 52.0% 31.2% 27.0%
Morio and Manning (2023) BERT-base 71.0% 63.5% 51.6% 73.6% 48.1% 37.2% 50.2% 31.9% 25.8%
Morio and Manning (2023) ClimateBERT 71.8% 64.0% 52.8% 74.4% 48.9% 39.0% 50.2% 32.2% 26.8%
Morio and Manning (2023) RoBERTa-base 71.6% 64.5% 53.1% 73.8% 49.6% 38.3% 50.4% 33.4% 26.6%
Morio and Manning (2023) Longformer-base 73.7% 66.9% 54.6% 75.9% 53.0% 40.8% 52.5% 36.1% 28.6%
Morio and Manning (2023) Longformer-large 73.9% 68.8% 57.3% 76.5% 55.0% 44.1% 53.6% 38.7% 31.5%
(a) Custom metrics on LobbyMap (Morio and Manning, 2023).
Table 20. Reported performance for the datasets on stance detection (Part 2)
C.7. Question Answering
Source embeddings Quest. Def. Concepts Gen. inf_3 inf_all
Schimanski et al. (2024b) random 0.037 0.037 0.037 0.037 0.037 0.037
Schimanski et al. (2024b) BM25 0.113 0.114 0.126 0.139 0.172 0.174
Schimanski et al. (2024b) ColBERTv2 0.109 0.094 0.112 0.137 0.130 0.137
Schimanski et al. (2024b) DRAGON+ 0.139 0.121 0.161 0.160 0.161 0.160
Schimanski et al. (2024b) GTE-base 0.161 0.153 0.171 0.154 0.171 0.174
Schimanski et al. (2024b) text-embedding-ada-002 0.163 0.143 0.152 0.163 0.176 0.176
Schimanski et al. (2024b) text-embedding-3-small 0.163 0.143 0.152 0.163 0.179 0.179
Schimanski et al. (2024b) text-embedding-3-large 0.167 0.143 0.152 0.163 0.179 0.179
(a) F1-score (custom) of the different embedding models and information retrieval approaches (question, definition, concepts, generic, inf_3, inf_all) (Schimanski et al., 2024b)
Test In-house annotation
Source Model F1 BLEU METEOR F1 BLEU METEOR
Rony et al. (2022) ALBERT 81.6% 0.678 0.808 66.1% 0.416 0.694
(b) F1-score, BLEU and METEOR on CCMRC (Rony et al., 2022)
Source Model F1
Luccioni et al. (2020) large 85.5%
Luccioni et al. (2020) base 82%
(c) F1-scores (macro average) of RoBERTa on question answering (Luccioni et al., 2020).
Source Model F1
Spokoyny et al. (2023) Majority 4.11%
Spokoyny et al. (2023) Random 12.14%
Spokoyny et al. (2023) SVM 86.00%
Spokoyny et al. (2023) BERT 84.57%
Spokoyny et al. (2023) RoBERTa 85.61%
Spokoyny et al. (2023) DistilRoBERTa 84.38%
Spokoyny et al. (2023) Longformer 84.35%
Spokoyny et al. (2023) SciBERT 84.43%
Spokoyny et al. (2023) ClimateBERT 84.80%
(d) F1-scores (macro average) on ClimaINS Spokoyny et al. (2023)
Source Model Question Answer
R@1 R@5 R@1 R@5
Thulke et al. (2024) bge-base-en-v1.5 54.8 71.5 81.8 92.1
Thulke et al. (2024) bge-large-en-v1.5 55.8 73.6 83.3 93.1
Thulke et al. (2024) gtr-t5-large 48.8 67.4 79.6 90.1
Thulke et al. (2024) gtr-t5-xxl 47.6 66.3 79.2 89.7
Thulke et al. (2024) gte-large 50.7 68.2 80.9 91.4
Thulke et al. (2024) ember-v1 49.5 68.6 79.7 91.1
Thulke et al. (2024) instructor-large 50.0 68.2 81.7 91.8
Thulke et al. (2024) instructor-xl 53.3 69.7 83.3 92.1
(e) Recall (R@1 and R@5) of retrieving the correct document given the question or the answer from the IFT dataset (Thulke et al., 2024).
Model CITIES STATES CORP
No Finetuning on CDP
BM25 0.055 0.084 0.153
MiniLM 0.099 0.120 0.320
Finetuned on CDP (In-Domain)
ClimateBERT 0.331 0.422 0.753
MiniLM 0.366 0.482 0.755
Best Model Finetuned on all
MiniLM 0.352 0.489 0.745
(f) MRR@10 scores on the three subsets of CLIMA-QA (Spokoyny et al., 2023). Models finetuned and evaluated on the same subset fall under In-Domain.
Task Raw Optimal Confidence
Net-zero target year 0.95 0.95 0.97
Reduction target year 0.84 0.90 0.90
Reduction base year 0.68 0.80 0.82
Reduction percentage 0.88 0.91 0.93
(g) Accuracy of the Q&A approach in assessing the quantitative properties of the net zero and reduction targets (Schimanski et al., 2023a)
Table 21. Reported performances for the datasets on Question/Answering and Retrieval tasks.
Source Model ClimaBench Pira 2.0 MCQ Exeter Misinf. Weight. Avg.
Acc. Acc. F1 (macro)
Thulke et al. (2024) Stability-3B 71.4% 48.7% 52.6% 62.8%
Thulke et al. (2024) Pythia-6.9B 63.6% 22.9% 29.9% 48.9%
Thulke et al. (2024) Falcon-7B 62.9% 19.8% 39.9% 48.3%
Thulke et al. (2024) Mistral-7B 73.1% 19.0% 63.7% 53.7%
Thulke et al. (2024) Llama-2-7B 68.5% 51.1% 59.4% 62.6%
Thulke et al. (2024) Jais-13B 66.9% 26.4% 54.2% 54.4%
Thulke et al. (2024) Jais-13B-Chat 65.8% 66.4% 61.3% 64.6%
Thulke et al. (2024) Llama-2-Chat-7B 67.8% 72.0% 64.3% 68.5%
Thulke et al. (2024) Llama-2-Chat-13B 68.6% 79.3% 68.6% 71.4%
Thulke et al. (2024) Llama-2-Chat-70B 72.7% 88.2% 72.5% 77.8%
Thulke et al. (2024) ClimateGPT-7B 75.3% 86.6% 65.9% 77.1%
Thulke et al. (2024) ClimateGPT-13B 75.0% 89.0% 70.0% 78.0%
Thulke et al. (2024) ClimateGPT-70B 72.4% 89.9% 72.5% 77.7%
Thulke et al. (2024) ClimateGPT-FSC-7B 59.3% 17.2% 45.1% 46.2%
Thulke et al. (2024) ClimateGPT-FSG-7B 53.1% 17.4% 41.5% 42.1%
(a) F1 scores (macro average) and Accuracy on multiple climate benchmarks (Thulke et al., 2024).
Source Model F1
Spokoyny et al. (2023) Majority 20.10%
Spokoyny et al. (2023) Random 24.09%
Spokoyny et al. (2023) SVM -%
Spokoyny et al. (2023) BERT 70.57%
Spokoyny et al. (2023) RoBERTa 71.14%
Spokoyny et al. (2023) DistilRoBERTa 69.27%
Spokoyny et al. (2023) Longformer 67.72%
Spokoyny et al. (2023) SciBERT 68.45%
Spokoyny et al. (2023) ClimateBERT 69.44%
(b) F1-scores (macro average) on ClimaBench (Spokoyny et al., 2023)
Source Model Acc.
Vaghefi et al. (2023) hybrid ChatClimate 4.38
Vaghefi et al. (2023) ChatClimate 4.15
Vaghefi et al. (2023) GPT-4 2.62
(c) Accuracy on a set of 10 questions, human annotations (Vaghefi et al., 2023).
Table 22. Reported performance for the datasets on Question/Answering (few-shot and RAG) Tasks.
C.8. Deceptive Technique
Source Model F1
Trained on Logic applied on LogicCLimate
Jin et al. (2022) Electra-SA 22.72%
Jin et al. (2022) Electra 27.23%
Trained on LogicClimate
Jin et al. (2022) Electra-SA 23.71%
Jin et al. (2022) Electra 29.37%
Baseline: Trained on Logic applied on Logic
Jin et al. (2022) Electra-SA 58.77%
Jin et al. (2022) Electra 53.31%
Jin et al. (2022) BERT 45.80%
(a) F1-scores (micro average) on fallacy detection (Jin et al., 2022). Electra-StructureAware is abbreviated Electra-SA
Source Model F1
Bhatia et al. (2021) BERT 59%
Bhatia et al. (2021) BERT* 62%
Bhatia et al. (2021) MTEXT 66%
Bhatia et al. (2021) MTEXT* 67%
Bhatia et al. (2021) MTEXTmulti 68%
Bhatia et al. (2021) MTEXT*multi 67%
Bhatia et al. (2021) SVM 49%
Bhatia et al. (2021) Human 70%
(b) F1-score (micro average) on neutralization techniques classification (Bhatia et al., 2021)
Source Model F1 (macro)
Coan et al. (2021) Logistic (unweighted) 68%
Coan et al. (2021) Logistic (weighted) 72%
Coan et al. (2021) SVM (unweighted) 66%
Coan et al. (2021) SVM (weighted) 72%
Coan et al. (2021) ULMFiT 72%
Coan et al. (2021) ULMFiT (weighted) 65%
Coan et al. (2021) ULMFiT (over sample) 55%
Coan et al. (2021) ULMFiT (focal Loss) 61%
Coan et al. (2021) ULMFiT-logistic 75%
Coan et al. (2021) ULMFiT-SVM 71%
Coan et al. (2021) RoBERTa 77%
Coan et al. (2021) RoBERTa-logistic 79%
Coan et al. (2021) Coder (Human) blAcc (macro): 87%
Thulke et al. (2024) Llama-2-Chat-70Bbla F1 (Bin.): 72.5%
Thulke et al. (2024) ClimateGPT-70Bblah F1 (Bin.): 72.5%
(c) F1-score (macro average and Binary) and Accuracy (macro average) on contrarian claim classification (Coan et al., 2021).
Table 23. Deceptive techniques
C.9. Environmental Performance Prediction
Source Model MSE
Lin et al. (2023b) Baseline 1.173
Lin et al. (2023b) Lasso 0.848
Lin et al. (2023b) Gradient Boosting 0.818
Lin et al. (2023b) NOCATE(RB) 0.776
Lin et al. (2023b) NOCATE(CB) 0.763
Lin et al. (2023b) NOCATE(DB) 0.756
Lin et al. (2023b) SUSTAINABLESIGNALS(RB) 0.762
Lin et al. (2023b) SUSTAINABLESIGNALS(CB) 0.753
Lin et al. (2023b) SUSTAINABLESIGNALS(DB) 0.736
(a) MSE on Finch Score prediction (Finch Score range between 0 and 10) (Lin et al., 2023b)
Source Model F1
Clarkson et al. (2020) Random Forest 89.48%
Clarkson et al. (2020) XGBoost 95.73%
(b) F1-score (macro average) on good/bad CSR performer classification (Clarkson et al., 2020).
Source Model RMSE wMAPE
Bronzini et al. (2023) OLS 7.76 7.9%
Uniform Random 40 66%
Normal Average 20 31%
(c) RMSE and wMAPE on ESG score prediction (ESG score range between 0 and 100) (Bronzini et al., 2023). We added the average RMSE and wMAPE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=50, scale=20) with the average value as prediction.
Source Model MAE RMSE
Sharma et al. (2024) CLIP+MLP 0.54 0.88
Sharma et al. (2024) Sent.-BERT+MLP 0.57 0.91
Uniform Random 3.97 4.88
Gaussian Average 1.59 1.99
(d) MAE and RMSE on ESG unmanaged risk score prediction (ESG risk between 0 and 13.5) (Sharma et al., 2024). We added the average RMSE and MAE for a uniform distribution of score with a random classifier; and for a normal distribution (mean=7, scale=2) with the average value a prediction.
Source Model Accuracy (Change) Accuracy (Direction of change)
Mehra et al. (2022) majority class 0.5791 0.5682
Mehra et al. (2022) BERT 0.5985 0.4317
Mehra et al. (2022) ESGBERT 0.6709 0.793
(e) Accuracy on ESG-risk change classification (Mehra et al., 2022)
Table 24. Environmental Performance Prediction